### v1

# Certain finite groups whose commuting conjugacy class graph satisfy Hansen-Vukičević conjecture

Abstract.
In this paper we compute first and second Zagreb indices of commuting conjugacy class graphs of finite groups and determine several classes of finite groups such that their commuting conjugacy class graphs satisfy Hansen-Vukičević conjecture.
Key words and phrases:
Commuting conjugacy class graph, Zagreb indices, finite group
2020 Mathematics Subject Classification:
20D60, 05C25, 05C09
*Corresponding author
1. Introduction
Let $\mathfrak{G}$ be the set of all graphs. A topological index is a function $T:\mathfrak{G}\to\mathbb{R}$ such that $T(\Gamma_{1})=T(\Gamma_{2})$ whenever the graphs $\Gamma_{1}$ and $\Gamma_{2}$ are isomorphic. Initially, topological indices were used to describe several chemical properties of molecules. Various topological indices have been defined since 1947 by using different parameters of graphs. Wiener index is the oldest topological index, introduced by Wiener [30], and it is a distance based index. Among the degree-based topological indices, the Zagreb indices are the first and were introduced by Gutman and Trinajstić [14] in 1972. They were used in examining the dependence of total $\pi$-electron energy on molecular structure. As noted in [22], Zagreb indices are also used in studying molecular complexity, chirality, ZE-isomerism and heterosystems etc. Later on, general mathematical properties of various topological indices are also studied by many mathematicians. A survey on mathematical properties of Zagreb indices can be found in [15].
Let $\Gamma$ be a simple undirected graph with vertex set $v(\Gamma)$ and edge set $e(\Gamma)$. The first and second Zagreb indices of $\Gamma$, denoted by $M_{1}(\Gamma)$ and $M_{2}(\Gamma)$ respectively, are defined as
$M_{1}(\Gamma)=\sum\limits_{v\in v(\Gamma)}\deg(v)^{2}\text{ and }M_{2}(\Gamma)=\sum\limits_{uv\in e(\Gamma)}\deg(u)\deg(v),$
where $\deg(v)$ is the number of edges incident on $v$ (called degree of $v$). Comparing first and second Zagreb indices, Hansen and Vukičević [16] posed the following conjecture in 2007.
Conjecture 1.1.
(Hansen-Vukičević Conjecture) For any simple finite graph $\Gamma$,
(1) $\dfrac{M_{2}(\Gamma)}{|e(\Gamma)|}\geq\dfrac{M_{1}(\Gamma)}{|v(\Gamma)|}.$
It was shown in [16] that the conjecture is not true if $\Gamma=K_{1,5}\sqcup K_{3}$. However, Hansen and Vukičević [16] showed that Conjecture 1.1 holds for chemical graphs. In [28], it was shown that the conjecture holds for trees with equality in (1) when $\Gamma$ is a star graph. In [18], it was shown that the conjecture holds for connected unicyclic graphs with equality when the graph is a cycle. The case when equality holds in (1) is studied extensively in [29]. A survey on comparing Zagreb indices can be found in [19]. However, the search of graphs validating or invalidating Conjecture 1.1 is not completed yet. Recently, Das et al. [8] have obtained various finite non-abelian groups such that their commuting/non-commuting graphs satisfy Hansen-Vukičević Conjecture. Das and Nath [9] also obtained several finite groups such that their super commuting graphs satisfy Hansen-Vukičević Conjecture. In this paper, we compute first and second Zagreb indices of commuting conjugacy class graphs of finite groups and determine several classes of finite groups such that their commuting conjugacy class graphs satisfy Hansen-Vukičević conjecture. In Section 2.1, we consider commuting conjugacy class graphs of dihedral groups $D_{2m}$ presented by $\langle a,b:a^{m}=b^{2}=1,bab^{-1}=a^{-1}\rangle$ for $m\geq 3$, dicyclic groups $Q_{4m}$ presented by $\langle a,b:a^{2m}=1,b^{2}=a^{n},bab^{-1}=a^{-1}\rangle$ for $m\geq 2$, semidihedral groups $SD_{8m}$ presented by $\langle a,b:a^{4m}=b^{2}=1,bab=a^{2m-1}\rangle$ for $m\geq 2$, the group $V_{8m}$ presented by $\langle a,b:a^{2m}=b^{4}=1,ba=b^{-1}a^{-1},b^{-1}a=a^{-1}b\rangle$ for $m\geq 1$, the group $U_{(n,m)}$ presented by $\langle a,b:a^{2n}=b^{m}=1,a^{-1}ba=b^{-1}\rangle$ for $m\geq 3$, $n\geq 2$ and the group $G(p,m,n)$ presented by $\langle a,b:a^{p^{m}}=b^{p^{m}}=[a,b]^{p}=1,[a,[a,b]]=[b,[a,b]]=1\rangle$ for any prime $p$ and $m,n\geq 1$. In Section 2.2, we consider commuting conjugacy class graphs of the groups $G$ when $G/Z(G)$ is isomorphic to $D_{2m}$, $\mathbb{Z}_{p}\times\mathbb{Z}_{p}$, Frobenious group of order $pq$ and $p^{2}q$ or any group of order $p^{3}$ for any two primes $p$ and $q$.
The commuting conjugacy class graph (in short CCC-graph) of $G$ is a simple undirected graph whose vertex set is the set of conjugacy classes of non-central elements of $G$ and two vertices $x^{G}$ (conjugacy class containing $x\in G$) and $y^{G}$ are adjacent if and only if there exists $x^{\prime}\in x^{G}$ and $y^{\prime}\in y^{G}$ such that $x^{\prime}y^{\prime}=y^{\prime}x^{\prime}$. We write $\mathcal{CCC}(G)$ to denote the CCC-graph of $G$. This graph was introduced by Herzog et al. [17] in 2009. In 2016, Mohammadian et al. [13] have characterized finite groups $G$ such that $\mathcal{CCC}(G)$ is triangle-free. In [26, 27], Salahshour and Ashrafi obtained structures of $\mathcal{CCC}(G)$ of various families of finite CA-groups. Salahshour also obtained $\mathcal{CCC}(G)$ for the groups whose central quotient is isomorphic to a dihedral group [25]. Recently, Rezaei et al. [23] obtained the structures of $\mathcal{CCC}(G)$ when the central quotient of $G$ is isomorphic to a Frobenious group of order $pq$ or $p^{2}q$ for any two primes $p$ and $q$. Various spectral aspects of $\mathcal{CCC}(G)$ have been considered in [4, 6, 11, 12]. In [5], genus of $\mathcal{CCC}(G)$ have been computed and determined several groups such that their CCC-graphs are planar, toroidal, double-toroidal or triple-toroidal. A survey on CCC-graphs of groups can be found in [7].
2. Zagreb indices of CCC-graphs
The following result, obtained in [8], is useful in our computations.
Theorem 2.1.
Let $\Gamma$ be the disjoint union of the graphs $\Gamma_{1},\Gamma_{2},\dots,\Gamma_{n}$. If $\Gamma_{i}=l_{i}K_{m_{i}}$ for $i=1,2,\dots,k$, where $K_{m_{i}}$’s are complete graphs on $m_{i}$ vertices and $l_{i}K_{m_{i}}$ is the disjoint union of $l_{i}$ copies of $K_{m_{i}}$, then
$M_{1}(\Gamma)=\sum_{i=1}^{k}l_{i}m_{i}(m_{i}-1)^{2}~{}~{}~{}\text{and}~{}~{}~{}M_{2}(\Gamma)=\sum_{i=1}^{k}l_{i}\dfrac{m_{i}(m_{i}-1)^{3}}{2}.$
2.1. Zagreb indices of CCC-graphs of certain well-known finite groups
In this section, we consider the CCC-graphs for the groups $D_{2m},Q_{4m},V_{8m},SD_{8m},U_{(n,m)}$ and $G(p,m,n)$ and compute their Zagreb indices to finally check if they satisfy Hansen-Vukičević conjecture. The structures of the CCC-graphs of the above mentioned groups are obtained in [27]. However, while obtaining the structure of CCC-graph of $U_{(n,m)}$ in [27] one case was missed which has been pointed out by Jannat and Nath in [10].
Theorem 2.2.
If $G$ is the dihedral group of order $2m$, where $m\geq 3$, then
$M_{1}(\mathcal{CCC}(G))=\begin{cases}\frac{(m-1)(m-3)^{2}}{8},&~{}~{}~{}~{}\text{ when }m\text{ is odd}\\ \frac{(m-2)(m-4)^{2}}{8},&~{}~{}~{}~{}\text{ when }m\text{ is even and }\frac{m}{2}\text{ is even}\\ \frac{(m-2)(m-4)^{2}}{8}+2,&~{}~{}~{}~{}\text{ when }m\text{ is even and }\frac{m}{2}\text{ is odd}\end{cases}$
and
$M_{2}(\mathcal{CCC}(G))=\begin{cases}\frac{(m-1)(m-3)^{3}}{32},&~{}~{}~{}~{}\text{ when }m\text{ is odd}\\ \frac{(m-2)(m-4)^{3}}{32},&~{}~{}~{}~{}\text{ when }m\text{ is even and }\frac{m}{2}\text{ is even}\\ \frac{(m-2)(m-4)^{3}}{32}+1,&~{}~{}~{}~{}\text{ when }m\text{ is even and }\frac{m}{2}\text{ is odd.}\end{cases}$
Further, $\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}\geq\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}$ with equality when $m=3,4$ and $6$.
Proof.
Case 1. $m$ is odd.
We have $\mathcal{CCC}(G)=K_{\frac{m-1}{2}}\sqcup K_{1}$. As such, $|v(\mathcal{CCC}(G))|=\frac{m-1}{2}+1=\frac{m+1}{2}$ and $|e(\mathcal{CCC}(G))|=\binom{\frac{m-1}{2}}{2}=\frac{(m-1)(m-3)}{8}$. Therefore, by Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))$ $=\frac{m-1}{2}\left(\frac{m-1}{2}-1\right)^{2}+0=\frac{(m-1)(m-3)^{2}}{8}$
and
$M_{2}(\mathcal{CCC}(G))$ $=\frac{\frac{m-1}{2}\left(\frac{m-1}{2}-1\right)^{3}}{2}+0=\frac{(m-1)(m-3)^{3}}{32}.$
Now
$\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{(m-1)(m-3)^{2}}{4(m+1)}\quad\text{ and }\quad\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}=\frac{(m-3)^{2}}{4}.$
Therefore
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}$ $=\frac{(m-3)^{2}}{4}\left(1-\frac{m-1}{m+1}\right)$
$=\frac{(m-3)^{2}}{4}\left(\frac{2}{m+1}\right)\geq 0$
for all $m\geq 3$.
Case 2. $m$ is even and $\frac{m}{2}$ is even.
We have $\mathcal{CCC}(G)=K_{\frac{m}{2}-1}\sqcup 2K_{1}$. As such, $|v(\mathcal{CCC}(G))|=\frac{m}{2}-1+2=\frac{m+2}{2}$ and $|e(\mathcal{CCC}(G))|=\binom{\frac{m}{2}-1}{2}=\frac{(m-2)(m-4)}{8}$. Therefore, by Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))$ $=\left(\frac{m}{2}-1\right)\left(\frac{m}{2}-1-1\right)^{2}+0=\frac{(m-2)(m-4)^{2}}{8}$
and
$M_{2}(\mathcal{CCC}(G))$ $=\frac{\left(\frac{m}{2}-1\right)\left(\frac{m}{2}-1-1\right)^{3}}{2}+0=\frac{(m-2)(m-4)^{3}}{32}.$
Now
$\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{(m-2)(m-4)^{2}}{4(m+2)}\quad\text{ and }\quad\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}=\frac{(m-4)^{2}}{4}.$
Therefore
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}$ $=\frac{(m-4)^{2}}{4}\left(1-\frac{m-2}{m+2}\right)$
$=\frac{(m-4)^{2}}{4}\left(\frac{4}{m+2}\right)\geq 0$
for all $m\geq 4$.
Case 3. $m$ is even and $\frac{m}{2}$ is odd.
We have $\mathcal{CCC}(G)=K_{\frac{m}{2}-1}\sqcup K_{2}$. As such, $|v(\mathcal{CCC}(G))|=\frac{m}{2}-1+2=\frac{m+2}{2}$ and $|e(\mathcal{CCC}(G))|=\binom{\frac{m}{2}-1}{2}+1=\frac{(m-2)(m-4)}{8}+1$. Therefore, by Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))$ $=\left(\frac{m}{2}-1\right)\left(\frac{m}{2}-1-1\right)^{2}+2(2-1)^{2}=\frac{(m-2)(m-4)^{2}}{8}+2$
and
$M_{2}(\mathcal{CCC}(G))$ $=\frac{\left(\frac{m}{2}-1\right)\left(\frac{m}{2}-1-1\right)^{3}}{2}+\frac{2(2-1)^{3}}{2}=\frac{(m-2)(m-4)^{3}}{32}+1.$
Therefore
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}$ $-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{\frac{(m-2)(m-4)^{3}}{32}+1}{\frac{(m-2)(m-4)}{8}+1}-\frac{\frac{(m-2)(m-4)^{2}}{8}+2}{\frac{m+2}{2}}$
$=\frac{1}{4}\left(\frac{4(m-2)(m-4)(m^{2}-10m+20)+32(m+2)-128}{(m+2)(m-2)(m-4)+8(m+2)}\right)$
$=\frac{m^{4}-16m^{3}+88m^{2}-200m+160+8(m-2)}{(m+2)(m-2)(m-4)+8(m+2)}$
$=\frac{m^{2}(m^{2}-16m+68)+20m(m-10)+8(m+18)}{(m+2)(m-2)(m-4)+8(m+2)}:=\frac{f(m)}{g(m)}.$
Note that $g(m)=16\times|e(\mathcal{CCC}(G))|\times|v(\mathcal{CCC}(G))|>0$. We have $f(6)=0$. Also, since $m(m-16)+68>0$ for all $m\geq 10$, therefore $\frac{f(m)}{g(m)}\geq 0$. Hence, the result follows. ∎
Theorem 2.3.
If $G$ is the dicyclic group of order $4m$, where $m\geq 2$, then
$M_{1}(\mathcal{CCC}(G))=\begin{cases}(m-1)(m-2)^{2},&\text{ when}m\text{ is even}\\ (m-1)(m-2)^{2}+2,&\text{ when }m\text{ is odd}\end{cases}$
and
$M_{2}(\mathcal{CCC}(G))=\begin{cases}\frac{(m-1)(m-2)^{3}}{2},&\text{ when }m\text{ is even}\\ \frac{(m-1)(m-2)^{3}}{2}+1,&\text{ when }m\text{ is odd.}\end{cases}$
Further, $\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}\geq\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}$ with equality when $m=2$ and $3$.
Proof.
Case 1. $m$ is even.
We have $\mathcal{CCC}(G)=K_{m-1}\sqcup 2K_{1}$. As such, $|v(\mathcal{CCC}(G))|=m-1+2=m+1$ and $|e(\mathcal{CCC}(G))|=\binom{m-1}{2}=\frac{(m-1)(m-2)}{2}$. By Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))=(m-1)(m-1-1)^{2}+0=(m-1)(m-2)^{2}$
and
$M_{2}(\mathcal{CCC}(G))=\frac{(m-1)(m-1-1)^{3}}{2}+0=\frac{(m-1)(m-2)^{3}}{2}.$
Therefore
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}$ $=\frac{\frac{(m-1)(m-2)^{3}}{2}}{\frac{(m-1)(m-2)}{2}}-\frac{(m-1)(m-2)^{2}}{m+1}$
$=(m-2)^{2}-\frac{(m-1)(m-2)^{2}}{m+1}$
$=\frac{(m-2)^{2}(m+1-m+1)}{m+1}=\frac{2(m-2)^{2}}{m+1}\geq 0$
for all $m\geq 2$.
Case 2. $m$ is odd.
We have $\mathcal{CCC}(G)=K_{m-1}\sqcup K_{2}$. As such, $|v(\mathcal{CCC}(G))|=m-1+2=m+1$ and $|e(\mathcal{CCC}(G))|=\binom{m-1}{2}+1=\frac{(m-1)(m-2)}{2}+1$. By Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))=(m-1)(m-1-1)^{2}+2(2-1)^{2}=(m-1)(m-2)^{2}+2$
and
$M_{2}(\mathcal{CCC}(G))=\frac{(m-1)(m-1-1)^{3}}{2}+\frac{2(2-1)^{3}}{2}=\frac{(m-1)(m-2)^{3}}{2}+1.$
Therefore
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}$ $=\frac{\frac{(m-1)(m-2)^{3}}{2}+1}{\frac{(m-1)(m-2)}{2}+1}-\frac{(m-1)(m-2)^{2}+2}{m+1}$
$=\frac{(m-1)(m-2)^{3}+2}{(m-1)(m-2)+2}-\frac{(m-1)(m-2)^{2}+2}{m+1}$
$=\frac{m^{4}-8m^{3}+22m^{2}-24m+9}{(m+1)(m-1)(m-2)+2(m+1)}$
$=\frac{m^{2}(m^{2}-8m+14)+8m(m-3)+9}{(m+1)(m-1)(m-2)+2(m+1)}:=\frac{f(m)}{g(m)}.$
Note that $g(m)=2\times|e(\mathcal{CCC}(G))|\times|v(\mathcal{CCC}(G))|>0$. We have $f(3)=0$ and $f(5)=64>0$. Since $m(m-8)+14>0$ for all $m\geq 7$, we have $\frac{f(m)}{g(m)}\geq 0$. Hence the result follows. ∎
Theorem 2.4.
If $G$ is the semidihedral group of order $8m$, where $m\geq 2$, then
$M_{1}(\mathcal{CCC}(G))=\begin{cases}(2m-1)(2m-2)^{2},&\text{ when }m\text{ is even}\\ (2m-2)(2m-3)^{2}+36,&\text{ when }m\text{ is odd}\end{cases}$
and
$M_{2}(\mathcal{CCC}(G))=\begin{cases}\frac{(2m-1)(2m-2)^{3}}{2},&\text{ when }m\text{ is even}\\ (m-1)(2m-3)^{3}+54,&\text{ when }m\text{ is odd.}\end{cases}$
Further, $\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}\geq\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}$ with equality when $m=3$.
Proof.
Case 1. $m$ is even.
We have $\mathcal{CCC}(G)=K_{2m-1}\sqcup 2K_{1}$. As such, $|v(\mathcal{CCC}(G))|=2m-1+2=2m+1$ and $|e(\mathcal{CCC}(G))|=\binom{2m-1}{2}+0=\frac{(2m-1)(2m-2)}{2}$. By Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))=(2m-1)(2m-1-1)^{2}+2(1-1)^{2}=(2m-1)(2m-2)^{2}$
and
$M_{2}(\mathcal{CCC}(G))=\frac{(2m-1)(2m-1-1)^{3}}{2}+\frac{2(1-1)^{3}}{2}=\frac{(2m-1)(2m-2)^{3}}{2}.$
Therefore
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}$ $=\frac{\frac{(2m-1)(2m-2)^{3}}{2}}{\frac{(2m-1)(2m-2)}{2}}-\frac{(2m-1)(2m-2)^{2}}{2m+1}$
$=(2m-2)^{2}-\frac{(2m-1)(2m-2)^{2}}{2m+1}$
$=\frac{(2m-2)^{2}(2m+1-2m+1)}{2m+1}=\frac{2(2m-2)^{2}}{2m+1}\geq 0$
for all $m\geq 1$.
Case 2. $m$ is odd.
We have $\mathcal{CCC}(G)=K_{2m-2}\sqcup K_{4}$. As such, $|v(\mathcal{CCC}(G))|=2m-2+4=2m+2$ and $|e(\mathcal{CCC}(G))|=\binom{2m-2}{2}+\binom{4}{2}=\frac{(2m-2)(2m-3)}{2}+6=(m-1)(2m-3)+6$. By Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))=(2m-2)(2m-2-1)^{2}+4(4-1)^{2}=(2m-2)(2m-3)^{2}+36$
and
$M_{2}(\mathcal{CCC}(G))=\frac{(2m-2)(2m-2-1)^{3}}{2}+\frac{4(4-1)^{3}}{2}=(m-1)(2m-3)^{3}+54.$
Therefore
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}$ $=\frac{(m-1)(2m-3)^{3}+54}{(m-1)(2m-3)+6}-\frac{(2m-2)(2m-3)^{2}+36}{2m+2}$
$=\frac{(m-1)(2m-3)^{3}+54}{(m-1)(2m-3)+6}-\frac{(m-1)(2m-3)^{2}+18}{m+1}$
$=\frac{8m^{4}-56m^{3}+120m^{2}-72m}{(m+1)(m-1)(2m-3)+6(m+1)}$
$=\frac{8m^{3}(m-7)+4m(30m-18)}{(m+1)(m-1)(2m-3)+6(m+1)}:=\frac{f(m)}{g(m)}.$
Note that $g(m)=\frac{1}{2}\times|e(\mathcal{CCC}(G))|\times|v(\mathcal{CCC}(G))|>0$. We have $f(3)=0$ and $f(5)=640>0$. Also, since $030m-18>$ 0 for all $m\geq 7$, therefore $\frac{f(m)}{g(m)}\geq 0$. Hence the result follows. ∎
Theorem 2.5.
If $G=V_{8m}$, where $m\geq 1$, then
$M_{1}(\mathcal{CCC}(G))=\begin{cases}(2m-2)(2m-3)^{2}+4,&\text{ when }m\text{ is even}\\ (2m-1)(2m-2)^{2},&\text{ when }m\text{ is odd}\end{cases}$
and
$M_{2}(\mathcal{CCC}(G))=\begin{cases}(m-1)(2m-3)^{3}+2,&\text{ when }m\text{ is even}\\ \frac{(2m-1)(2m-2)^{3}}{2}+1,&\text{ when }m\text{ is odd.}\end{cases}$
Further, $\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}\geq\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}$ with equality when $m=1$ and $2$.
Proof.
Case 1. $m$ is even.
We have $\mathcal{CCC}(G)=K_{2m-2}\sqcup 2K_{2}$. As such, $|v(\mathcal{CCC}(G))|=2m-2+4=2m+2$ and $|e(\mathcal{CCC}(G))|=\binom{2m-2}{2}+2=\frac{(2m-2)(2m-3)}{2}+2=(m-1)(2m-3)+2$. By Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))=(2m-2)(2m-2-1)^{2}+2\times 2(2-1)^{2}=(2m-2)(2m-3)^{2}+4$
and
$M_{2}(\mathcal{CCC}(G))=\frac{(2m-2)(2m-2-1)^{3}}{2}+2\times\frac{2(2-1)^{3}}{2}=(m-1)(2m-3)^{3}+2.$
Therefore
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}$ $=\frac{(m-1)(2m-3)^{3}+2}{(m-1)(2m-3)+2}-\frac{(2m-2)(2m-3)^{2}+4}{2m+2}$
$=\frac{(m-1)(2m-3)^{3}+2}{(m-1)(2m-3)+2}-\frac{(m-1)(2m-3)^{2}+2}{m+1}$
$=\frac{8m^{4}-48m^{3}+104m^{2}-96m+32}{(m+1)(m-1)(2m-3)+2(m+1)}$
$=\frac{8m^{3}(m-6)+8m(13m-12)+32}{(m+1)(m-1)(2m-3)+2(m+1)}:=\frac{f(m)}{g(m)}.$
Note that $g(m)=\frac{1}{2}\times|e(\mathcal{CCC}(G))|\times|v(\mathcal{CCC}(G))|>0$. We have $f(2)=0$ and $f(4)=288>0$. Also, since $013m-12>$ 0 for all $m\geq 6$ we have $\frac{f(m)}{g(m)}>0$.
Case 2. $m$ is odd.
We have $\mathcal{CCC}(G)=K_{2m-1}\sqcup 2K_{1}$. Therefore, the result follows from the first case of Theorem 2.4.
This completes the proof. ∎
Theorem 2.6.
If $G=U_{(n,m)}$, where $m\geq 3$ and $n\geq 2$, then
$M_{1}(\mathcal{CCC}(G))=\begin{cases}\frac{1}{8}\left(16n(n-1)^{2}+(mn-2n)(mn-2n-2)^{2}\right),\\ \hskip 170.71652pt\text{ when }m\text{ is even and }\frac{m}{2}\text{ is even}\\ \frac{1}{8}\left(16n(2n-1)^{2}+(mn-2n)(mn-2n-2)^{2}\right),\\ \hskip 170.71652pt\text{ when }m\text{ is even and }\frac{m}{2}\text{ is odd}\\ \frac{1}{8}\left(8n(n-1)^{2}+(mn-n)(mn-n-2)^{2}\right),\text{ when }m\text{ isodd}\end{cases}$
and
$M_{2}(\mathcal{CCC}(G))=\begin{cases}\frac{1}{32}\left(32n(n-1)^{3}+(mn-2n)(mn-2n-2)^{3}\right),\\ \hskip 170.71652pt\text{ when }m\text{ is even and }\frac{m}{2}\text{ is even}\\ \frac{1}{32}\left(32n(2n-1)^{3}+(mn-2n)(mn-2n-2)^{3}\right),\\ \hskip 170.71652pt\text{ when }m\text{ is even and }\frac{m}{2}\text{ is odd}\\ \frac{1}{32}\left(16n(n-1)^{3}+(mn-n)(mn-n-2)^{3}\right),\text{ when }m\text{ is odd.}\end{cases}$
Further, $\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}\geq\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}$ with equality when $m=3,4$ and $6$.
Proof.
Case 1. $m$ is even and $\frac{m}{2}$ is even.
We have $\mathcal{CCC}(G)=2K_{n}\sqcup K_{n(\frac{m}{2}-1)}$. As such, $|v(\mathcal{CCC}(G))|=2n+n(\frac{m}{2}-1)=\frac{1}{2}(mn+2n)$ and $|e(\mathcal{CCC}(G))|=2\binom{n}{2}+\binom{\frac{mn}{2}-n}{2}=\frac{1}{8}(8n(n-1)+(mn-2n)(mn-2n-2))$. By Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))$ $=2n(n-1)^{2}+\left(\frac{mn}{2}-n\right)\left(\frac{mn}{2}-n-1\right)^{2}$
$=\frac{1}{8}\left(16n(n-1)^{2}+(mn-2n)(mn-2n-2)^{2}\right)$
and
$M_{2}(\mathcal{CCC}(G))$ $=2n\frac{(n-1)^{3}}{2}+\left(\frac{mn}{2}-n\right)\frac{\left(\frac{mn}{2}-n-1\right)^{3}}{2}+$
$=\frac{1}{32}\left(32n(n-1)^{3}+(mn-2n)(mn-2n-2)^{3}\right).$
Therefore
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{\frac{1}{32}\left(32n(n-1)^{3}+(mn-2n)(mn-2n-2)^{3}\right)}{\frac{1}{8}(8n(n-1)+(mn-2n)(mn-2n-2))}$
$\hskip 156.49014pt-\frac{\frac{1}{8}\left(16n(n-1)^{2}+(mn-2n)(mn-2n-2)^{2}\right)}{\frac{1}{2}(mn+2n)}$
$=\frac{1}{4}\times\frac{4n^{4}(m-2)(m-4)^{2}(mn-4)}{8n(n-1)(mn+2n)+(mn+2n)(mn-2n)(mn-2n-2)}:=\frac{1}{4}\times\frac{f_{1}(m,n)}{g_{1}(m,n)}.$
Note that $g_{1}(m,n)=16\times|e(\mathcal{CCC}(G))|\times|v(\mathcal{CCC}(G))|>0$. We have $f_{1}(4,n)=0$. Also, since $m-2>0$, $m-4>0$ and $mn-4>0$ for all $m\geq 6$ and $n\geq 2$, therefore, $\frac{1}{4}\times\frac{f_{1}(m,n)}{g_{1}(m,n)}\geq 0$. Hence the result follows.
Case 2. $m$ is even and $\frac{m}{2}$ is odd.
We have $\mathcal{CCC}(G)=K_{2n}\sqcup K_{n(\frac{m}{2}-1)}$. As such, $|v(\mathcal{CCC}(G))|=2n+n(\frac{m}{2}-1)=\frac{1}{2}(mn+2n)$ and $|e(\mathcal{CCC}(G))|=\binom{2n}{2}+\binom{\frac{mn}{2}-n}{2}=\frac{1}{8}(8n(2n-1)+(mn-2n)(mn-2n-2))$. By Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))$ $=2n(2n-1)^{2}+\left(\frac{mn}{2}-n\right)\left(\frac{mn}{2}-n-1\right)^{2}$
$=\frac{1}{8}\left(16n(2n-1)^{2}+(mn-2n)(mn-2n-2)^{2}\right)$
and
$M_{2}(\mathcal{CCC}(G))$ $=2n\frac{(2n-1)^{3}}{2}+\left(\frac{mn}{2}-n\right)\frac{\left(\frac{mn}{2}-n-1\right)^{3}}{2}$
$=\frac{1}{32}\left(32n(2n-1)^{3}+(mn-2n)(mn-2n-2)^{3}\right).$
Therefore
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{\frac{1}{32}\left(32n(2n-1)^{3}+(mn-2n)(mn-2n-2)^{3}\right)}{\frac{1}{8}(8n(2n-1)+(mn-2n)(mn-2n-2))}$
$\hskip 156.49014pt-\frac{\frac{1}{8}\left(16n(2n-1)^{2}+(mn-2n)(mn-2n-2)^{2}\right)}{\frac{1}{2}(mn+2n)}$
$=\frac{1}{4}\times\frac{4n^{4}(m-2)(m-6)^{2}(mn+2n-4)}{8n(2n-1)(mn+2n)+(mn+2n)(mn-2n)(mn-2n-2)}:=\frac{1}{4}\times\frac{f_{2}(m,n)}{g_{2}(m,n)}.$
Note that $g_{2}(m,n)=16\times|e(\mathcal{CCC}(G))|\times|v(\mathcal{CCC}(G))|>0$. We have $f_{2}(6,n)=0$. Also, since $m-2>0$, $m-6>0$ and $mn+2n-4>0$ for all $m\geq 10$ and $n\geq 2$, therefore, $\frac{1}{4}\times\frac{f_{2}(m,n)}{g_{2}(m,n)}\geq 0$. Hence the result follows.
Case 3. $m$ is odd.
We have $\mathcal{CCC}(G)=K_{n}\sqcup K_{\frac{n(m-1)}{2}}$. As such, $|v(\mathcal{CCC}(G))|=\frac{mn-n}{2}+n=\frac{1}{2}(mn+n)$ and $|e(\mathcal{CCC}(G))|=\binom{\frac{mn-n}{2}}{2}+\binom{n}{2}=\frac{1}{8}((mn-n)(mn-n-2)+4n(n-1)$. By Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))$ $=\frac{(mn-n)}{2}\left(\frac{mn-n}{2}-1\right)^{2}+n(n-1)^{2}$
$=\frac{1}{8}\left((mn-n)(mn-n-2)^{2}+8n(n-1)^{2}\right)$
and
$M_{2}(\mathcal{CCC}(G))$ $=\frac{(mn-n)}{2}\frac{\left(\frac{mn-n}{2}-1\right)^{3}}{2}+n\frac{(n-1)^{3}}{2}$
$=\frac{1}{32}\left((mn-n)(mn-n-2)^{3}+16n(n-1)^{3}\right).$
Therefore
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{\frac{1}{32}\left((mn-n)(mn-n-2)^{3}+16n(n-1)^{3}\right)}{\frac{1}{8}((mn-n)(mn-n-2)+4n(n-1))}$
$\hskip 156.49014pt-\frac{\frac{1}{8}\left((mn-n)(mn-n-2)^{2}+8n(n-1)^{2}\right)}{\frac{1}{2}(mn+n)}$
$=\frac{1}{4}\times\frac{2m^{3}n^{4}(mn-6n-4)+16m^{2}n^{5}+4mn^{4}(14m-30)+6n^{5}(2m-3)+72n^{4}}{(mn+n)(mn-n)(mn-n-2)+4n(n-1)(mn+n)}$
$:=\frac{1}{4}\times\frac{f_{3}(m,n)}{g_{3}(m,n)}.$
Note that $g_{3}(m,n)=16\times|e(\mathcal{CCC}(G))|\times|v(\mathcal{CCC}(G))|>0$. We have $f_{3}(3,n)=0$, $f_{3}(5,n)=192n^{5}-128n^{4}>0$, $f_{3}(7,n)=1536n^{5}-768n^{4}>0$ and $f_{3}(9,n)=5760n^{5}-2304n^{4}>0$ for all $n\geq 1$. Also, since $n(m-6)-4=mn-6n-4>0$, $014m-30>$ 0 and $02m-3>$ 0 for all $m\geq 11$ and $n\geq 1$, therefore, $\frac{1}{4}\times\frac{f_{3}(m,n)}{g_{3}(m,n)}\geq 0$. Hence the result follows. ∎
We conclude this section with the following result.
Theorem 2.7.
If $G=G(p,m,n)$, where $p$ is a prime and $m,n\geq 1$, then
$M_{1}(\mathcal{CCC}(G))$ $=2p^{3m+3n-3}-6p^{3m+3n-4}+6p^{3m+3n-5}-2p^{3m+3n-6}-4p^{2m+2n-2}+p^{3m+n-4}$
$\quad-4p^{2m+2n-4}+p^{3m+n}-4p^{3m+n-1}+6p^{3m+n-2}-4p^{3m+n-3}+8p^{2m+2n-3}$
$\quad-2p^{2m+n}+6p^{2m+n-1}-6p^{2m+n-2}+2p^{2m+n-3}+p^{m+n}-p^{m+n-2}$
and
$M_{2}(\mathcal{CCC}(G))$ $=\frac{1}{2}\left(2p^{4m+4n-4}-8p^{4m+4n-5}+6p^{3m+3n-6}-8p^{4m+4n-7}+2p^{4m+4n-8}-p^{m+n}\right.$
$\quad\left.-6p^{3m+3n-3}+18p^{3m+3n-4}-18p^{3m+3n-5}+12p^{4m+4n-6}+6p^{2m+2n-2}\right.$
$\quad\left.+6p^{2m+2n-4}+p^{4m+n}-5p^{4m+n-1}+10p^{4m+n-2}-10p^{4m+n-3}+5p^{4m+n-4}\right.$
$\quad\left.+12p^{3m+n-1}-18p^{3m+n-2}+12p^{3m+n-3}-3p^{3m+n-4}+3p^{2m+n}-9p^{2m+n-1}\right.$
$\quad\left.-12p^{2m+2n-3}-3p^{3m+n}-p^{4m+n-5}+9p^{2m+n-2}-3p^{2m+n-3}+p^{m+n-2}\right).$
Further, $\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}\geq\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}$ with equality when $n=1$.
Proof.
We have $\mathcal{CCC}(G)=K_{p^{m-1}(p^{n}-p^{n-1})}\sqcup K_{p^{n-1}(p^{m}-p^{m-1})}\sqcup(p^{n}-p^{n-1})K_{p^{m-n}(p^{n}-p^{n-1})}$
$=2K_{(p^{m+n-1}-p^{m+n-2})}\sqcup(p^{n}-p^{n-1})K_{(p^{m}-p^{m-1})}$. As such, $|v(\mathcal{CCC}(G))|=2(p^{m+n-1}-p^{m+n-2})+(p^{n}-p^{n-1})(p^{m}-p^{m-1})=p^{m+n}-p^{m+n-2}$ and $|e(\mathcal{CCC}(G))|=2\binom{p^{m+n-1}-p^{m+n-2}}{2}+(p^{n}-p^{n-1})\binom{p^{m}-p^{m-1}}{2}$ $=\frac{1}{2}\left(2p^{2m+2n-2}-4p^{2m+2n-3}+2p^{2m+2n-4}+p^{2m+n}-3p^{2m+n-1}\right.\\ \left.+3p^{2m+n-2}-p^{2m+n-3}-p^{m+n}+p^{m+n-2}\right)$. By Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))=2(p^{m+n-1}-p^{m+n-2})$ $(p^{m+n-1}-p^{m+n-2}-1)^{2}$
$+(p^{n}-p^{n-1})(p^{m}-p^{m-1})(p^{m}-p^{m-1}-1)^{2}$
and
$M_{2}(\mathcal{CCC}(G))=2(p^{m+n-1}-p^{m+n-2})$$\times\frac{(p^{m+n-1}-p^{m+n-2}-1)^{3}}{2}$
$+(p^{n}-p^{n-1})(p^{m}-p^{m-1})\times\frac{(p^{m}-p^{m-1}-1)^{3}}{2}.$
Hence, we get the required expressions for $M_{1}(\mathcal{CCC}(G))$ and $M_{2}(\mathcal{CCC}(G))$ on simplification.
We have $\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{f(p,m,n)}{g(p,m,n)}$, where
$f(p,m,n)$ $=2p^{5m+5n-4}-12p^{5m+5n-5}+30p^{5m+5n-6}-40p^{5m+5n-7}+30p^{5m+5n-8}$
$\quad-12p^{5m+5n-9}+2p^{5m+5n-10}-2p^{5m+4n-3}+12p^{5m+4n-4}-30p^{5m+4n-5}$
$\quad+40p^{5m+4n-6}-30p^{5m+4n-7}+12p^{5m+4n-8}-2p^{5m+4n-9}-4p^{4m+4n-3}$
$\quad+20p^{4m+4n-4}-40p^{4m+4n-5}+40p^{4m+4n-6}-20p^{4m+4n-7}+4p^{4m+4n-8}$
$\quad-2p^{5m+3n-2}+12p^{5m+3n-3}-30p^{5m+3n-4}+40p^{5m+3n-5}-30p^{5m+3n-6}$
$\quad+12p^{5m+3n-7}-2p^{5m+3n-8}+8p^{4m+3n-2}-40p^{4m+3n-3}+80p^{4m+3n-4}$
$\quad-80p^{4m+3n-5}+40p^{4m+3n-6}-8p^{4m+3n-7}+2p^{5m+2n-1}-12p^{5m+2n-2}$
$\quad+30p^{5m+2n-3}-40p^{5m+2n-4}+30p^{5m+2n-5}-12p^{5m+2n-6}+2p^{5m+2n-7}$
$\quad-4p^{4m+2n-1}+20p^{4m+2n-2}-40p^{4m+2n-3}+40p^{4m+2n-4}-20p^{4m+2n-5}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+4p^{4m+2n-6}.$
and
$g(p,m,n)$ $=(p^{m+n}-p^{m+n-2})(2p^{2m+2n-2}-4p^{2m+2n-3}+2p^{2m+2n-4}+p^{2m+n}-3p^{2m+n-1}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+3p^{2m+n-2}-p^{2m+n-3}-p^{m+n}+p^{m+n-2}).$
On factorization, we have $f(p,m,n)=2p^{4m+2n-10}(p-1)^{5}(p-p^{n})^{2}(p^{m}(p^{n+1}-p^{n}-p)+p^{2}(p^{m}-2))$. Clearly, $f(p,m,1)=0$. For all prime $p\geq 2,m\geq 1$ and $n>1$, we have $p^{n+1}-p^{n}-p>0$ and $p^{m}-2\geq 0$. As such, $f(p,m,n)>0$ and $g(p,m,n)=2\times|e(\mathcal{CCC}(G))|\times|v(\mathcal{CCC}(G))|>0$. Thus, $\frac{f(p,m,n)}{g(p,m,n)}\geq 0$ and the result follows. ∎
2.2. Zagreb indices of CCC-graphs of groups with a given central quotient
In this section, we consider the CCC-graphs of the groups $G$ when $G/Z(G)$ is isomorphic to $D_{2m}$, $\mathbb{Z}_{p}\times\mathbb{Z}_{p}$, Frobenious group of order $pq$ and $p^{2}q$ or any group of order $p^{3}$ for any two primes $p$ and $q$. The structures of CCC-graphs of groups when its central quotient is isomorphic to a dihedral group have been obtained in [25, Theorem 1.2]. Using which we get the following result.
Theorem 2.8.
Let $G$ be a finite group such that $\frac{G}{Z(G)}\cong D_{2m}$, where $m\geq 3.$ Then
$M_{1}(\mathcal{CCC}(G))=\begin{cases}\frac{1}{8}\left((mx-x)(mx-x-2)^{2}+2x(x-2)^{2}\right),&\text{ when }m\text{ is even}\\ \frac{1}{8}\left((mx-x)(mx-x-2)^{2}+8x(x-1)^{2}\right),&\text{ when }m\text{ is odd}\end{cases}$
and
$M_{2}(\mathcal{CCC}(G))=\begin{cases}\frac{1}{32}\left((mx-x)(mx-x-2)^{3}+2x(x-2)^{3}\right),&\text{ when }m\text{ is even}\\ \frac{1}{32}\left((mx-x)(mx-x-2)^{3}+16x(x-1)^{3}\right),&\text{ when }m\text{is odd,}\end{cases}$
where $x=|Z(G)|$. Further, $\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}>\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}$.
Proof.
Case 1. $m$ is even.
We have $\mathcal{CCC}(G)=K_{\frac{x(m-1)}{2}}\sqcup 2K_{\frac{x}{2}}$. As such, $|v(\mathcal{CCC}(G))|=\frac{x(m-1)}{2}+2\times\frac{x}{2}=\frac{x(m+1)}{2}$ and $|e(\mathcal{CCC}(G))|=\binom{\frac{mx-x}{2}}{2}+2\binom{\frac{x}{2}}{2}=\frac{1}{8}((mx-x)(mx-x-2)+2x(x-2))$. By Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))$ $=\frac{(mx-x)}{2}\left(\frac{mx-x}{2}-1\right)^{2}+2\times\frac{x}{2}\left(\frac{x}{2}-1\right)^{2}$
$=\frac{1}{8}\left((mx-x)(mx-x-2)^{2}+2x(x-2)^{2}\right)$
and
$M_{2}(\mathcal{CCC}(G))$ $=\frac{(mx-x)}{2}\frac{\left(\frac{mx-x}{2}-1\right)^{3}}{2}+2\times\frac{x}{2}\frac{\left(\frac{x}{2}-1\right)^{3}}{2}$
$=\frac{1}{32}\left((mx-x)(mx-x-2)^{3}+2x(x-2)^{3}\right).$
Therefore
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{\frac{1}{32}\left((mx-x)(mx-x-2)^{3}+2x(x-2)^{3}\right)}{\frac{1}{8}((mx-x)(mx-x-2)+2x(x-2))}$
$\hskip 156.49014pt-\frac{\frac{1}{8}\left((mx-x)(mx-x-2)^{2}+2x(x-2)^{2}\right)}{\frac{1}{2}x(m+1)}$
$=\frac{1}{4}\times\frac{(mx-x)(mx-x-2)^{3}+2x(x-2)^{3}}{(mx-x)(mx-x-2)+2n(n-2)}-\frac{(mx-x)(mx-x-2)^{2}+2x(x-2)^{2}}{mx+x}$
$=\frac{1}{4}\times\frac{m^{3}x^{3}-4m^{2}x^{3}+4mx^{3}-4m^{2}x^{2}+16mx^{2}-16x^{2}}{mx(mx+x)(mx-2x-2)}$
$=\frac{1}{4}\times\frac{x^{2}(m-2)^{2}(mx-4)}{mx(mx+x)(mx-2x-2)}>0$
for all $m\geq 4$.
Case 2. $m$ is odd.
We have $\mathcal{CCC}(G)=K_{\frac{x(m-1)}{2}}\sqcup K_{x}$. As such, $|v(\mathcal{CCC}(G))|=\frac{mx-x}{2}+x=\frac{1}{2}(mx+x)$ and $|e(\mathcal{CCC}(G))|=\binom{\frac{mx-x}{2}}{2}+\binom{x}{2}=\frac{1}{8}((mx-x)(mx-x-2)+4x(x-1))$. By Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))$ $=\frac{(mx-x)}{2}\left(\frac{mx-x}{2}-1\right)^{2}+x(x-1)^{2}$
$=\frac{1}{8}\left((mx-x)(mx-x-2)^{2}+8x(x-1)^{2}\right)$
and
$M_{2}(\mathcal{CCC}(G))$ $=\frac{(mx-x)}{2}\frac{\left(\frac{mx-x}{2}-1\right)^{3}}{2}+x\frac{(x-1)^{3}}{2}$
$=\frac{1}{32}\left((mx-x)(mx-x-2)^{3}+16x(x-1)^{3}\right).$
Therefore
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{\frac{1}{32}\left((mx-x)(mx-x-2)^{3}+16x(x-1)^{3}\right)}{\frac{1}{8}((mx-x)(mx-x-2)+4x(x-1))}$
$\hskip 156.49014pt-\frac{\frac{1}{8}\left((mx-x)(mx-x-2)^{2}+8x(x-1)^{2}\right)}{\frac{1}{2}(mx+x)}$
$=\frac{1}{4}\times\frac{2m^{3}x^{4}(mx-6x-4)+16m^{2}x^{5}+4mx^{4}(14m-30)+6x^{5}(2m-3)+72x^{4}}{(mx+x)(mx-x)(mx-x-2)+4x(x-1)(mx+x)}$
$:=\frac{1}{4}\frac{f(m,x)}{g(m,x)}.$
Note that $g(m,x)=16\times|e(\mathcal{CCC}(G))|\times|v(\mathcal{CCC}(G))|>0$. We have $f(3,x)=0$, $f(5,x)=192x^{5}-128x^{4}>0$, $f(7,x)=1536x^{5}-768x^{4}>0$ and $f(9,x)=5760x^{5}-2304x^{4}>0$ for all $x\geq 1$. Also, since $x(m-6)-4=mx-6x-4>0$, $014m-30>$ 0 and $02m-3>$ 0 for all $m\geq 11$ and $x\geq 1$, we have $\frac{1}{4}\frac{f(m,x)}{g(m,x)}\geq 0$. Hence the result follows. ∎
The structures of CCC-graphs of groups when its central quotient is isomorphic to $\mathbb{Z}_{p}\times\mathbb{Z}_{p}$ (for any prime $p$) have been obtained in [26, Theotem 3.1]. Therefore, we get the following result.
Theorem 2.9.
Let $G$ be a finite group and $\frac{G}{Z(G)}\cong\mathbb{Z}_{p}\times\mathbb{Z}_{p}$, where $p$ is a prime. Then
$M_{1}(\mathcal{CCC}(G))=\frac{x(p^{2}-1)}{p^{3}}\left(p^{2}x^{2}-2px^{2}+x^{2}-2p^{2}x+2px+p^{2}\right)\text{ \quad and }$
$M_{2}$ $(\mathcal{CCC}(G))$
$=\frac{x(p^{2}-1)}{2p^{4}}\left(p^{3}x^{3}-3p^{2}x^{3}+3px^{3}-x^{3}-3p^{3}x^{2}+6p^{2}x^{2}-3px^{2}\right.\left.+3p^{3}x-3p^{2}x-p^{3}\right),$
where $x=|Z(G)|$. Further, $\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}=\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}$.
Proof.
We have $\mathcal{CCC}(G)=(p+1)K_{\frac{x(p-1)}{p}}$. As such, $|v(\mathcal{CCC}(G))|=(p+1)\times\frac{x(p-1)}{p}=\frac{x(p^{2}-1)}{p}$ and $|e(\mathcal{CCC}(G))|=(p+1)\binom{\frac{x(p-1)}{p}}{2}=\frac{x(p^{2}-1)}{2p^{2}}(px-x-p)$. By Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))$ $=(p+1)\frac{x(p-1)}{p}\left(\frac{x(p-1)}{p}-1\right)^{2}$
$=\frac{x(p^{2}-1)}{p^{3}}\left(p^{2}x^{2}-2px^{2}+x^{2}-2p^{2}x+2px+p^{2}\right)$
and
$M_{2}$ $(\mathcal{CCC}(G))=(p+1)\frac{x(p-1)}{p}\frac{\left(\frac{x(p-1)}{p}-1\right)^{3}}{2}$
$=\frac{x(p^{2}-1)}{2p^{4}}\left(p^{3}x^{3}-3p^{2}x^{3}+3px^{3}-x^{3}-3p^{3}x^{2}+6p^{2}x^{2}-3px^{2}+3p^{3}x-3p^{2}x-p^{3}\right).$
Now
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}$ $=\frac{p^{3}x^{3}-3p^{2}x^{3}+3px^{3}-x^{3}-3p^{3}x^{2}+6p^{2}x^{2}-3px^{2}+3p^{3}x-3p^{2}x-p^{3}}{p^{2}(px-x-p)}$
$=\frac{(px-x-p)(p^{2}x^{2}-2px^{2}+x^{2}-2p^{2}x+2px+p^{2})}{p^{2}(px-x-p)}$
$=\frac{p^{2}x^{2}-2px^{2}+x^{2}-2p^{2}x+2px+p^{2}}{p^{2}}=\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}.$
Hence, the result follows. ∎
Corollary 2.10.
If $G$ is a group of order $p^{n}$ and $|Z(G)|=p^{n-2}$, where $p$ is a prime and $n\geq 3$, then
$M_{1}(\mathcal{CCC}(G))$
$=p^{3n-5}-2p^{3n-6}+2p^{3n-8}-p^{3n-9}-2p^{2n-3}+2p^{2n-4}+2p^{2n-5}-2p^{2n-6}+p^{n-1}-p^{n-3}$
and
$M_{2}(\mathcal{CCC}(G))$ $=p^{4n-7}-3p^{4n-8}+2p^{4n-9}+2p^{4n-10}-3p^{4n-11}+p^{4n-12}-3p^{3n-5}+6p^{3n-6}$
$\quad\quad-6p^{3n-8}+3p^{3n-9}+3p^{2n-3}-3p^{2n-4}-3p^{2n-5}+3p^{2n-6}-p^{n-1}+p^{n-3}.$
Further, $\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}=\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}$.
Proof.
Clearly, $\frac{G}{Z(G)}\cong\mathbb{Z}_{p}\times\mathbb{Z}_{p}$ and $\frac{(p-1)|Z(G)|}{p}=(p-1)p^{n-3}$. Therefore, the result follows from Theorem 2.9. ∎
Putting $n=3$ in Corollary 2.10 we get the following result.
Corollary 2.11.
If $G$ is a non-abelian group of order $p^{3}$ then
$M_{1}(\mathcal{CCC}(G))=p^{4}-4p^{3}+3p^{2}+4p-4,\,M_{2}(\mathcal{CCC}(G))=p^{5}-6p^{4}+2p^{3}+9p^{3}-2p^{2}-12p+8$
and $\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}=\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}$.
Now we consider finite groups whose central quotient is isomorphic to certain Frobenious groups. The CCC-graphs of these groups have been obtained in [23].
Theorem 2.12.
Let $p$ and $q$ be two primes. Let $G$ be a finite group such that $\frac{G}{Z(G)}\cong\mathbb{Z}_{q}\rtimes_{f}\mathbb{Z}_{p}$, the Frobenious group of order $pq$ with kernel $\mathbb{Z}_{q}$ and complement $\mathbb{Z}_{p}$. Then
$M_{1}(\mathcal{CCC}(G))$ $=\frac{x}{p^{3}}\left(q^{3}x^{2}-3q^{2}x^{2}+3qx^{2}+p^{6}x^{2}-3p^{5}x^{2}+3p^{4}x^{2}-p^{3}x^{2}-x^{2}-2pq^{2}x\right.$
$\quad\quad\quad\left.+4pqx-2p^{5}x+4p^{4}x-2p^{3}x-2px+p^{2}q+p^{4}-p^{3}-p^{2}\right)\text{ \quad and }$
$M_{2}(\mathcal{CCC}(G))$ $=\frac{x}{2p^{4}}\left(q^{4}x^{3}-4q^{3}x^{3}+6q^{2}x^{3}-4px^{3}-p^{3}q-4p^{7}x^{3}+6p^{6}x^{3}-4p^{5}x^{3}+3px^{2}\right.$
$\quad\left.+3p^{2}q^{2}x-3pq^{3}x^{2}+9pq^{2}x^{2}-9pqx^{2}-3p^{7}x^{2}+9p^{6}x^{2}-9p^{5}x^{2}+3p^{4}x^{2}\right.$
$\quad\left.+p^{4}x^{3}-6p^{2}qx+3p^{6}x-6p^{5}x+3p^{4}x+3p^{2}x+x^{3}+p^{8}x^{3}-p^{5}+p^{4}+p^{3}\right)$
where $x=|Z(G)|$. Further, $\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}\geq\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}$.
Proof.
We have $\mathcal{CCC}(G)=K_{\frac{x(q-1)}{p}}\sqcup K_{x(p-1)}$, where $x=|Z(G)|$. As such, $|v(\mathcal{CCC}(G))|=\frac{x(q-1)}{p}+x(p-1)=\frac{x}{p}(p^{2}+q-p-1)$ and $|e(\mathcal{CCC}(G))|=\binom{\frac{x(q-1)}{p}}{2}+\binom{x(p-1)}{2}=\frac{x}{2p^{2}}(p^{2}x+q^{2}x+p^{4}x+x-2qx-pq-2p^{3}x-p^{3}+p^{2}+p)$. Also, by Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))$ $=\frac{x(q-1)}{p}\left(\frac{x(q-1)}{p}-1\right)^{2}+x(p-1)\left(x(p-1)-1\right)^{2}$
$=\frac{x}{p^{3}}\left(q^{3}x^{2}-3q^{2}x^{2}+3qx^{2}+p^{6}x^{2}-3p^{5}x^{2}+3p^{4}x^{2}-p^{3}x^{2}-x^{2}-2pq^{2}x\right.$
$\quad\quad\quad\left.+4pqx-2p^{5}x+4p^{4}x-2p^{3}x-2px+p^{2}q+p^{4}-p^{3}-p^{2}\right)\text{ \quad and }$
$M_{2}(\mathcal{CCC}(G))$ $=\frac{x(q-1)}{p}\frac{\left(\frac{x(q-1)}{p}-1\right)^{3}}{2}+x(p-1)\frac{\left(x(p-1)-1\right)^{3}}{2}$
$=\frac{x}{2p^{4}}\left(q^{4}x^{3}-4q^{3}x^{3}+6q^{2}x^{3}-4px^{3}-p^{3}q-4p^{7}x^{3}+6p^{6}x^{3}-4p^{5}x^{3}+3px^{2}\right.$
$\quad\left.+3p^{2}q^{2}x-3pq^{3}x^{2}+9pq^{2}x^{2}-9pqx^{2}-3p^{7}x^{2}+9p^{6}x^{2}-9p^{5}x^{2}+3p^{4}x^{2}\right.$
$\,\,\,\left.+p^{4}x^{3}-6p^{2}qx+3p^{6}x-6p^{5}x+3p^{4}x+3p^{2}x+x^{3}+p^{8}x^{3}-p^{5}+p^{4}+p^{3}\right).$
Therefore
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}$
$\qquad\qquad=\frac{f(p,q,x)}{p^{2}(p^{2}+q-p-1)(p^{2}x+q^{2}x+p^{4}x+x-2qx-pq-2p^{3}x-p^{3}+p^{2}+p)}$
$\qquad\qquad:=\frac{f(p,q,x)}{g(p,q,x)},$
where
$f(p,q,x)$ $=p^{2}q^{4}x^{3}-pq^{4}x^{3}-p^{4}q^{3}x^{3}+2p^{3}q^{3}x^{3}-5p^{2}q^{3}x^{3}+4pq^{3}x^{3}-p^{6}q^{2}x^{3}+3p^{5}q^{2}x^{3}$
$\quad-5p^{3}q^{2}x^{3}+9p^{2}q^{2}x^{3}-6pq^{2}x^{3}+p^{8}qx^{3}-4p^{7}qx^{3}+8p^{6}qx^{3}-10p^{5}qx^{3}-p^{3}x^{3}$
$\quad+4p^{4}qx^{3}+4p^{3}qx^{3}-7p^{2}qx^{3}+4pqx^{3}-p^{8}x^{3}+4p^{7}x^{3}-7p^{6}x^{3}+7p^{5}x^{3}-px^{3}$
$\quad+2p^{2}x^{3}-2p^{3}q^{3}x^{2}+2p^{2}q^{3}x^{2}+4p^{5}q^{2}x^{2}-8p^{4}q^{2}x^{2}-6p^{2}q^{2}x^{2}-2p^{7}qx^{2}$
$\quad+6p^{6}qx^{2}-14p^{5}qx^{2}+18p^{4}qx^{2}-14p^{3}qx^{2}+6p^{2}qx^{2}+2p^{7}x^{2}-6p^{6}x^{2}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+10p^{5}x^{2}-3p^{4}x^{3}-10p^{4}x^{2}+6p^{3}x^{2}-2p^{2}x^{2}.$
On factorization, we have $f(p,q,x)=px^{2}(p-1)(q-1)(p+q-p^{2}-1)^{2}(p^{2}x+qx-x-2p-px)$. Suppose that $p^{2}x+qx-x-2p-px=p(x(p-1)-2)+x(q-1):=S_{1}+S_{2}$, where $S_{1}=p(x(p-1)-2)$ and $S_{2}=x(q-1)>0$ for all prime $q$ and $x\geq 1$. Clearly, $S_{1}\geq 0$ for all prime $p\geq 3$ and $x\geq 1$. Now, $f(2,q,1)=2(q-1)(q-3)^{2}(q-3)\geq 0$ for all $q\geq 3$. Also, $g(p,q,x)=\frac{2p^{5}}{x^{2}}\times|e(\mathcal{CCC}(G))|\times|v(\mathcal{CCC}(G))|>0$. Therefore, $\frac{f(p,q,x)}{g(p,q,x)}\geq 0$ and hence the result follows. ∎
Theorem 2.13.
Let $G$ be a finite group such that $\frac{G}{Z(G)}$ is isomorphic to a Frobenious group of order $p^{2}q$, where $p$ and $q$ are primes. Then $\mathcal{CCC}(G)$ satisfies Hansen-Vukičević conjecture if $\frac{G}{Z(G)}\cong A_{4}$ or $p<q$ and $|\frac{G}{Z(G)}|\neq 12$ or $p>q$.
Proof.
We have the following three cases.
Case 1. $\frac{G}{Z(G)}\cong A_{4}$.
We have $\mathcal{CCC}(G)=K_{2x}\sqcup K_{x}$ or $K_{2x}\sqcup K_{\frac{x}{2}}$, where $x=|Z(G)|$. If $\mathcal{CCC}(G)=K_{2x}\sqcup K_{x}$ then, by Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))$ $=2x(2x-1)^{2}+x(x-1)^{2}=9x^{3}-10x^{2}+3x\text{ \quad and}$
$M_{2}(\mathcal{CCC}(G))$ $=2x\frac{(2x-1)^{3}}{2}+x\frac{(x-1)^{3}}{2}=\frac{1}{2}\left(17x^{4}-27x^{3}+15x^{2}-3x\right).$
Also, $|v(\mathcal{CCC}(G))|=2x+x=3x$ and $|e(\mathcal{CCC}(G))|=\binom{2x}{2}+\binom{x}{2}=\frac{1}{2}(5x^{2}-3x)$. Therefore, for all $x\geq 1$, we have
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{6x^{5}-4x^{4}}{3x(5x^{2}-3x)}>0.$
If $\mathcal{CCC}(G)=K_{2x}\sqcup K_{\frac{x}{2}}$ then, by Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))$ $=2x(2x-1)^{2}+\frac{x}{2}\left(\frac{x}{2}-1\right)^{2}=\frac{1}{8}\left(65x^{3}-68x^{2}+20x\right)\text{ \quad and}$
$M_{2}(\mathcal{CCC}(G))$ $=2x\frac{(2x-1)^{3}}{2}+\frac{x}{2}\frac{(\frac{x}{2}-1)^{3}}{2}=\frac{1}{32}\left(257x^{4}-390x^{3}+204x^{2}-40x\right).$
Also, $|v(\mathcal{CCC}(G))|=2x+\frac{x}{2}=\frac{5x}{2}$ and $|e(\mathcal{CCC}(G))|=\binom{2x}{2}+\binom{\frac{x}{2}}{2}=\frac{1}{8}(17x^{2}-10x)$. Therefore, for all $x\geq 1$, we have
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{180x^{5}-144x^{4}}{20x(17x^{2}-10x)}>0.$
Case 2. $p<q$ and $|\frac{G}{Z(G)}|\neq 12$.
We have $\mathcal{CCC}(G)=K_{\frac{x(q-1)}{p^{2}}}\sqcup K_{x(p^{2}-1)}$, where $x=|Z(G)|$. As such, $|v(\mathcal{CCC}(G))|=\frac{x(q-1)}{p^{2}}+x(p^{2}-1)=\frac{1}{p^{2}}(p^{4}x+qx-p^{2}x-x)$ and $|e(\mathcal{CCC}(G))|=\binom{\frac{x(q-1)}{p^{2}}}{2}+\binom{x(p^{2}-1)}{2}=\frac{1}{2p^{4}}(q^{2}x^{2}-2qx^{2}+p^{8}x^{2}-2p^{6}x^{2}+p^{4}x^{2}+x^{2}-p^{2}qx-p^{6}x+p^{4}x+p^{2}x)$. By Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))$ $=\frac{x(q-1)}{p^{2}}\left(\frac{x(q-1)}{p^{2}}-1\right)^{2}+x(p^{2}-1)\left(x(p^{2}-1)-1\right)^{2}$
$=\frac{1}{p^{6}}\left(q^{3}x^{3}-3q^{2}x^{3}+3qx^{3}+p^{12}x^{3}-3p^{10}x^{3}+3p^{8}x^{3}-p^{6}x^{3}-2p^{2}q^{2}x^{2}\right.$
$\quad\quad\quad\quad\quad\quad\quad\quad\left.-x^{3}+4p^{2}qx^{2}+4p^{8}x^{2}-2p^{10}x^{2}-2p^{6}x^{2}+p^{4}qx\right.$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.-2p^{2}x^{2}+p^{8}x-p^{4}x-p^{6}x\right)\text{ \quad and }$
$M_{2}(\mathcal{CCC}(G))$ $=\frac{x(q-1)}{p^{2}}\frac{\left(\frac{x(q-1)}{p^{2}}-1\right)^{3}}{2}+x(p^{2}-1)\frac{\left(x(p^{2}-1)-1\right)^{3}}{2}$
$=\frac{1}{2p^{8}}\left(q^{4}x^{4}-4q^{3}x^{4}+6q^{2}x^{4}-4qx^{4}+p^{16}x^{4}-4p^{14}x^{4}-4p^{10}x^{4}+6p^{12}x^{4}\right.$
$\quad\quad\quad\quad\left.+x^{4}+p^{8}x^{4}-3p^{2}q^{3}x^{3}+9p^{2}q^{2}x^{3}-9p^{2}qx^{3}+9p^{12}x^{3}-3p^{14}x^{3}\right.$
$\quad\quad\quad\quad\quad\left.+3p^{8}x^{3}-9p^{10}x^{3}+3p^{4}q^{2}x^{2}+3p^{2}x^{3}-6p^{4}qx^{2}+3p^{12}x^{2}\right.$
$\quad\quad\quad\quad\quad\left.-6p^{10}x^{2}+3p^{4}x^{2}+3p^{8}x^{2}-p^{10}x-p^{6}qx+p^{6}x+p^{8}x\right).$
Therefore, $\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{f_{1}(p,q,x)}{g_{1}(p,q,x)}$, where $f_{1}(p,q,x)=p^{2}x^{4}(p-1)(p+1)(p^{4}-p^{2}-q+1)^{2}(q-1)(p^{2}(p^{2}x-x-2)+x(q-1))$ after factorization and $g_{1}(p,q,x)=2p^{10}\times|e(\mathcal{CCC}(G))|\times|v(\mathcal{CCC}(G))|>0$. For all prime $p\geq 2$ and $x\geq 1$, we have $x(p^{2}-1)>2$. As such, $f_{1}(p,q,x)>0$ and therefore, $\frac{f_{1}(p,q,x)}{g_{1}(p,q,x)}>0$.
Case 3. $p>q$.
We have $\mathcal{CCC}(G)=K_{x(q-1)}\sqcup K_{\frac{x(p^{2}-1)}{q}}$ or $K_{x(q-1)}\sqcup(p+1)K_{\frac{x(p-1)}{pq}}$. If $\mathcal{CCC}(G)=K_{x(q-1)}\sqcup K_{\frac{x(p^{2}-1)}{q}}$ then, by Theorem 2.1, we have
$M_{1}$ $(\mathcal{CCC}(G))=x(q-1)\left(x(q-1)-1\right)^{2}+\frac{x(p^{2}-1)}{q}\left(\frac{x(p^{2}-1)}{q}-1\right)^{2}$
$\qquad=\frac{1}{q^{3}}\left(q^{6}x^{3}-3q^{5}x^{3}+3q^{4}x^{3}-q^{3}x^{3}+p^{6}x^{3}-3p^{4}x^{3}+3p^{2}x^{3}-2q^{5}x^{2}-x^{3}\right.$
$\qquad\quad\quad\quad\left.+4q^{4}x^{2}-2q^{3}x^{2}-2p^{4}qx^{2}+4p^{2}qx^{2}-2qx^{2}+q^{4}x-q^{3}x+p^{2}q^{2}x-q^{2}x\right)$
and
$M_{2}(\mathcal{CCC}(G))$ $=x(q-1)\frac{\left(x(q-1)-1\right)^{3}}{2}+\frac{x(p^{2}-1)}{q}\frac{\left(\frac{x(p^{2}-1)}{q}-1\right)^{3}}{2}$
$=\frac{1}{2q^{4}}\left(q^{8}x^{4}-4q^{7}x^{4}+6q^{6}x^{4}-4q^{5}x^{4}+p^{8}x^{4}+q^{4}x^{4}-4p^{6}x^{4}+6p^{4}x^{4}\right.$
$\quad\quad\quad\left.-4p^{2}x^{4}+x^{4}-3q^{7}x^{3}+9q^{6}x^{3}-9q^{5}x^{3}+3q^{4}x^{3}-3p^{6}qx^{3}+9p^{4}qx^{3}\right.$
$\quad\quad\quad\left.-9p^{2}qx^{3}+3q^{6}x^{2}+3qx^{3}-6q^{5}x^{2}+3p^{4}q^{2}x^{2}+3q^{4}x^{2}-6p^{2}q^{2}x^{2}\right.$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.+3q^{2}x^{2}-q^{5}x+q^{4}x+q^{3}x-p^{2}q^{3}x\right).$
Also, $|v(\mathcal{CCC}(G))|=x(q-1)+\frac{x(p^{2}-1)}{q}=\frac{1}{q}(q^{2}x-qx+p^{2}x-x)$ and $|e(\mathcal{CCC}(G))|=\binom{x(q-1)}{2}+\binom{\frac{x(p^{2}-1)}{q}}{2}=\frac{1}{2q^{2}}(q^{4}x^{2}-2q^{3}x^{2}+q^{2}x^{2}+p^{4}x^{2}-2p^{2}x^{2}-q^{3}x+x^{2}-p^{2}qx+q^{2}x+qx)$. Therefore, $\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{f_{2}(p,q,x)}{g_{2}(p,q,x)}$, where $f_{2}(p,q,x)=qx^{4}(p-1)(p+1)(q-1)(q^{2}-p^{2}-q+1)^{2}(x(p^{2}-q-1)+q(qx-2))$ after factorization and $g_{2}(p,q,x)=2q^{5}\times|e(\mathcal{CCC}(G))|\times|v(\mathcal{CCC}(G))|>0$. Since $p>q$, so $p^{2}-q>1$ and $qx>2$ for all prime $q\geq 2$ and $x\geq 1$. As such, $f_{2}(p,q,x)>0$ and therefore, $\frac{f_{2}(p,q,x)}{g_{2}(p,q,x)}>0$.
If $\mathcal{CCC}(G)=K_{x(q-1)}\sqcup(p+1)K_{\frac{x(p-1)}{pq}}$ then, by Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))$ $=x(q-1)\left(x(q-1)-1\right)^{2}+(p+1)\frac{x(p-1)}{pq}\left(\frac{x(p-1)}{pq}-1\right)^{2}$
$=\frac{1}{p^{3}q^{3}}\left(p^{3}q^{6}x^{3}-3p^{3}q^{5}x^{3}+3p^{3}q^{4}x^{3}-p^{3}q^{3}x^{3}+p^{4}x^{3}-2p^{3}x^{3}+2px^{3}\right.$
$\quad\quad\quad\left.-2p^{3}q^{5}x^{2}-x^{3}+4p^{3}q^{4}x^{2}-2p^{4}qx^{2}-2p^{3}q^{3}x^{2}+2p^{2}qx^{2}\right.$
$\quad\quad\quad\left.+2p^{3}qx^{2}+p^{3}q^{4}x-2pqx^{2}+p^{4}q^{2}x-p^{3}q^{3}x-p^{2}q^{2}x\right)$
and
$M_{2}(\mathcal{CCC}(G))$ $=x(q-1)\frac{\left(x(q-1)-1\right)^{3}}{2}+(p+1)\frac{x(p-1)}{pq}\frac{\left(\frac{x(p-1)}{pq}-1\right)^{3}}{2}$
$=\frac{1}{2p^{4}q^{4}}\left(p^{4}q^{8}x^{4}-4p^{4}q^{7}x^{4}+6p^{4}q^{6}x^{4}-4p^{4}q^{5}x^{4}+p^{4}q^{4}x^{4}+p^{5}x^{4}-3p^{4}x^{4}\right.$
$\quad\quad\quad\quad\left.+2p^{2}x^{4}+2p^{3}x^{4}-3px^{4}+x^{4}-3p^{4}q^{7}x^{3}+9p^{4}q^{6}x^{3}-9p^{4}q^{5}x^{3}\right.$
$\quad\quad\quad\quad\left.+3p^{4}q^{4}x^{3}-3p^{5}qx^{3}+6p^{4}qx^{3}-6p^{2}qx^{3}+3p^{4}q^{6}x^{2}+3pqx^{3}\right.$
$\quad\quad\quad\quad\left.-6p^{4}q^{5}x^{2}+3p^{5}q^{2}x^{2}+3p^{4}q^{4}x^{2}-3p^{3}q^{2}x^{2}-3p^{4}q^{2}x^{2}+3p^{2}q^{2}x^{2}\right.$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.-p^{4}q^{5}x+p^{4}q^{4}x+p^{3}q^{3}x-p^{5}q^{3}x\right).$
Also, $|v(\mathcal{CCC}(G))|=x(q-1)+(p+1)\frac{x(p-1)}{pq}=\frac{1}{pq}(pq^{2}x-pqx+p^{2}x-x)$ and $|e(\mathcal{CCC}(G))|=\binom{x(q-1)}{2}+(p+1)\binom{\frac{x(p-1)}{pq}}{2}=\frac{1}{2p^{2}q^{2}}(p^{2}q^{4}x^{2}-2p^{2}q^{3}x^{2}+p^{2}q^{2}x^{2}+p^{3}x^{2}-px^{2}-p^{2}x^{2}+x^{2}-p^{2}q^{3}x+p^{2}q^{2}x+pqx-p^{3}qx)$. Therefore, $\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{f_{3}(p,q,x)}{g_{3}(p,q,x)}$, where $f_{3}(p,q,x)=pqx^{4}(p-1)(p+1)(q-1)(pq^{2}-pq-p+1)^{2}(pq(qx-x-2)+x(p-1))$ after factorization and $g_{3}(p,q,x)=2p^{5}q^{5}\times|e(\mathcal{CCC}(G))|\times|v(\mathcal{CCC}(G))|>0$. Since $x\geq p>q$, so $x(q-1)>2$. As such, $f_{3}(p,q,x)>0$ and therefore, $\frac{f_{3}(p,q,x)}{g_{3}(p,q,x)}>0$.
In all the cases, we have $\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}>\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}$ and hence the result follows.
Theorem 2.14.
Let $G$ be a non-abelian group such that $\frac{G}{Z(G)}$ is abelian and $|\frac{G}{Z(G)}|=p^{3}$, where $p$ is a prime. Then $\mathcal{CCC}(G)$ satisfies Hansen-Vukičević conjecture.
Proof.
We have $\mathcal{CCC}(G)=K_{m}\sqcup p^{2}K_{n}$ or $(p^{2}+p+1)K_{n}$, where $m=\frac{x(p^{2}-1)}{p},n=\frac{x(p-1)}{p^{2}}$ and $x=|Z(G)|$. If $\mathcal{CCC}(G)=K_{m}\sqcup p^{2}K_{n}$ then, by Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))$ $=\frac{x(p^{2}-1)}{p}\left(\frac{x(p^{2}-1)}{p}-1\right)^{2}+p^{2}\times\frac{x(p-1)}{p^{2}}\left(\frac{x(p-1)}{p^{2}}-1\right)^{2}$
$=\frac{x(p-1)}{p^{4}}\left(p^{6}x^{2}-2p^{4}x^{2}+2p^{2}x^{2}-2p^{5}x+2p^{4}+p^{5}x^{2}-2p^{3}x^{2}-px^{2}\right.$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.-2p^{4}x+4p^{2}x+p^{3}+x^{2}\right)\quad\text{ and }$
$M_{2}(\mathcal{CCC}(G))$ $=\frac{x(p^{2}-1)}{p}\frac{\left(\frac{x(p^{2}-1)}{p}-1\right)^{3}}{2}+p^{2}\times\frac{x(p-1)}{p^{2}}\frac{\left(\frac{x(p-1)}{p^{2}}-1\right)^{3}}{2}$
$=\frac{x(p-1)}{2p^{6}}\left(p^{9}x^{3}+p^{8}x^{3}-3p^{8}x^{2}-3p^{7}x^{3}-3p^{7}x^{2}-3p^{6}x^{3}+6p^{6}x^{2}\right.$
$\quad\quad\quad\quad\quad\quad\left.+3p^{7}x+3p^{6}x+3p^{5}x^{3}+6p^{5}x^{2}+3p^{4}x^{3}-6p^{4}x^{2}-6p^{4}x\right.$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.+3p^{3}x^{2}-4p^{2}x^{3}-p^{5}-2p^{6}-x^{3}-3p^{2}x^{2}+3px^{3}\right).$
Also, $|v(\mathcal{CCC}(G))|=\frac{x(p^{2}-1)}{p}+p^{2}\times\frac{x(p-1)}{p^{2}}=x(p-1)(p+2)$ and $|e(\mathcal{CCC}(G))|=\binom{\frac{x(p^{2}-1)}{p}}{2}+p^{2}\binom{\frac{x(p-1)}{p^{2}}}{2}=\frac{x(p-1)}{2p^{2}}(p^{3}x+p^{2}x-2p^{2}-2x-p)$. Therefore
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}$
$\,=\frac{1}{g(p,x)}\left(p^{10}x^{3}+2p^{9}x^{3}-3p^{9}x^{2}-3p^{8}x^{3}-5p^{8}x^{2}-8p^{7}x^{3}+3p^{7}x+3p^{6}x^{3}+17p^{6}x^{2}\right.$
$\quad\quad\quad\quad\quad\left.+11p^{5}x^{3}-9p^{5}x+3p^{8}x-15p^{4}x^{2}-8p^{3}x^{3}-p^{6}-2p^{7}+3px^{3}+3p^{3}x^{2}\right.$
$\quad\quad\quad\quad\quad\left.-2p^{2}x^{3}+p^{4}x^{3}+3p^{2}x^{2}+7p^{7}x^{2}-8p^{5}x^{2}-3p^{6}x+6p^{3}x+2p^{5}+p^{4}+px^{2}\right)$
$\,=\frac{1}{g(p,x)}\left(p(p^{2}-1)\left(p^{4}x^{3}(p^{3}-2p-4)+p^{4}x^{2}(p^{2}x-2x-5p)+x^{2}(2px-3p-3x-1)\right.\right.$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.\left.+p^{6}x^{2}(x-3)+3p^{5}x+p^{2}x(5x^{2}-4x-6)+p^{4}(4x^{2}+3x-2)\right.\right.$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.\left.+p^{3}(x^{3}+12x^{2}-1)\right)\right)$
$\,:=\frac{f(p,x)}{g(p,x)},$
where $g(p,x)=p^{4}(p+2)(p^{3}x+p^{2}x-2p^{2}-2x-p)=p^{4}(p+2)(x(p^{3}-2)-p+p^{2}(x-2))>0$ since $p^{3}-2>p$ and $x>2$. Now, $f(2,x)=750x^{3}-1290x^{2}+720x-240>0$ for all $x\geq p^{2}\geq 4$. Also, for all $p>2$, we have $p^{3}-2p-4>0,p^{2}x-2x-5p>0,2px-3p-3x-1>0$ and $05x^{2}-4x-6>$ 0. Thus, $\frac{f(p,x)}{g(p,x)}>0$.
If $\mathcal{CCC}(G)=(p^{2}+p+1)K_{n}$ then, by Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))$ $=(p^{2}+p+1)\frac{x(p-1)}{p^{2}}\left(\frac{x(p-1)}{p^{2}}-1\right)^{2}$
$=\frac{x(p^{3}-1)}{p^{6}}\left(p^{2}x^{2}-2px^{2}+x^{2}-2p^{3}x+2p^{2}x+p^{4}\right)\text{ \quad and }$
$M_{2}(\mathcal{CCC}(G))=(p^{2}+p+1)\frac{x(p-1)}{p^{2}}\frac{\left(\frac{x(p-1)}{p^{2}}-1\right)^{3}}{2}$
$=\frac{x(p^{3}-1)}{2p^{8}}\left(p^{3}x^{3}-3p^{2}x^{3}+3px^{3}-x^{3}-3p^{4}x^{2}+6p^{3}x^{2}-3p^{2}x^{2}+3p^{5}x-3p^{4}x-p^{6}\right).$
Also, $|v(\mathcal{CCC}(G))|=(p^{2}+p+1)\frac{x(p-1)}{p^{2}}=\frac{x(p^{3}-1)}{p^{2}}$ and $|e(\mathcal{CCC}(G))|=(p^{2}+p+1)\binom{\frac{x(p-1)}{p^{2}}}{2}$ $=\frac{x(p^{3}-1)}{2p^{4}}(px-x-p^{2})$. Therefore
$\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{p^{2}x^{2}-2px^{2}+x^{2}-2p^{3}x+2p^{2}x+p^{4}}{p^{4}}$
and
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}$ $=\frac{p^{3}x^{3}-3p^{2}x^{3}+3px^{3}-x^{3}-3p^{4}x^{2}+6p^{3}x^{2}-3p^{2}x^{2}+3p^{5}x-3p^{4}x-p^{6}}{p^{4}(px-x-p^{2})}$
$=\frac{(px-x-p^{2})(p^{2}x^{2}-2px^{2}+x^{2}-2p^{3}x+2p^{2}x+p^{4})}{p^{4}(px-x-p^{2})}$
$=\frac{p^{2}x^{2}-2px^{2}+x^{2}-2p^{3}x+2p^{2}x+p^{4}}{p^{4}}=\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}.$
In both the cases, we have $\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}\geq\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}$ and hence the result follows. ∎
Corollary 2.15.
Let $G$ be a non-abelian group of order $p^{n}$ such that $\frac{G}{Z(G)}$ is abelian and $|Z(G)|=p^{n-3}$, where $p$ is a prime and $n\geq 4$. Then $\mathcal{CCC}(G)$ satisfies Hansen-Vukičević conjecture.
Proof.
Clearly $|\frac{G}{Z(G)}|=p^{3}$ and so the result follows from Theorem 2.14. ∎
In the remaining part of this section we consider non-abelian groups $G$ such that $\frac{G}{Z(G)}$ is a group of order $p^{3}$ for any prime $p$. The CCC-graphs of these groups have been obtained in [26].
Theorem 2.16.
Let $G$ be a non-abelian group such that $\frac{G}{Z(G)}$ is non-abelian and $|\frac{G}{Z(G)}|=p^{3}$, where $p$ is a prime. Then $\mathcal{CCC}(G)$ satisfies Hansen-Vukičević conjecture.
Proof.
We have $\mathcal{CCC}(G)=K_{m}\sqcup kpK_{n_{1}}\sqcup(p-k)K_{n_{2}},(kp+1)K_{n_{1}}\sqcup(p+1-k)K_{n_{2}},K_{m}\sqcup pK_{n_{2}},(p^{2}+p+1)K_{n_{1}}$ or $K_{n_{1}}\sqcup(p+1)K_{n_{2}}$, where $x=|Z(G)|,m=\frac{x(p^{2}-1)}{p},n_{1}=\frac{x(p-1)}{p^{2}},n_{2}=\frac{x(p-1)}{p}$ and $1\leq k\leq p$. If $\mathcal{CCC}(G)=K_{m}\sqcup kpK_{n_{1}}\sqcup(p-k)K_{n_{2}}$ then, by Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))$ $=\frac{x(p^{2}-1)}{p}\left(\frac{x(p^{2}-1)}{p}-1\right)^{2}+kp\frac{x(p-1)}{p^{2}}\left(\frac{x(p-1)}{p^{2}}-1\right)^{2}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+(p-k)\frac{x(p-1)}{p}\left(\frac{x(p-1)}{p}-1\right)^{2}$
$=\frac{x(p-1)}{p^{6}}\left(p^{8}x^{2}+p^{7}x^{2}-2p^{7}x-p^{6}x^{2}-4p^{6}x-4p^{5}x^{2}+2p^{4}x^{2}+p^{3}x^{2}\right.$
$\quad\quad\quad\quad\quad\quad\left.+2p^{4}x+p^{5}+4p^{5}x+2p^{6}-2kp^{2}x^{2}+kpx^{2}-4kp^{4}x\right.$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.+2kp^{3}x-kp^{5}x^{2}+2kp^{4}x^{2}+2kp^{5}x\right)\text{ \quad and}$
$M_{2}(\mathcal{CCC}(G))$ $=\frac{x(p^{2}-1)}{p}\frac{\left(\frac{x(p^{2}-1)}{p}-1\right)^{3}}{2}+kp\frac{x(p-1)}{p^{2}}\frac{\left(\frac{x(p-1)}{p^{2}}-1\right)^{3}}{2}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+(p-k)\frac{x(p-1)}{p}\frac{\left(\frac{x(p-1)}{p}-1\right)^{3}}{2}$
$=\frac{x(p-1)}{2p^{8}}\left(p^{11}x^{3}-3p^{9}x^{3}-2p^{5}x^{3}-3p^{10}x^{2}+3p^{8}x^{2}-6p^{6}x^{2}+3p^{9}x\right.$
$\quad\quad\quad\quad\quad\quad\left.-6p^{7}x-2p^{8}+p^{10}x^{3}-2p^{8}x^{3}+6p^{6}x^{3}-p^{4}x^{3}-3p^{9}x^{2}\right.$
$\quad\quad\quad\quad\quad\quad\left.+12p^{7}x^{2}-3p^{5}x^{2}+6p^{8}x-3p^{6}x-p^{7}+2kp^{4}x^{3}-3kp^{3}x^{3}\right.$
$\quad\quad\quad\quad\quad\quad\left.+3kp^{2}x^{3}-kpx^{3}+6kp^{4}x^{2}-3kp^{3}x^{2}+6kp^{6}x-3kp^{5}x\right.$
$\quad\quad\quad\quad\quad\quad\left.-kp^{7}x^{3}+3kp^{6}x^{3}-3kp^{5}x^{3}+3kp^{7}x^{2}-6kp^{6}x^{2}-3kp^{7}x\right).$
Also, $|v(\mathcal{CCC}(G))|=\frac{x(p^{2}-1)}{p}+kp\frac{x(p-1)}{p^{2}}+(p-k)\frac{x(p-1)}{p}=\frac{x(p-1)}{p^{2}}(2p^{2}+p)$ and $|e(\mathcal{CCC}(G))|=\binom{\frac{x(p^{2}-1)}{p}}{2}+kp\binom{\frac{x(p-1)}{p^{2}}}{2}+(p-k)\binom{\frac{x(p-1)}{p^{2}}}{2}=\frac{x(p-1)}{2p^{4}}(p^{5}x+2p^{4}x-2p^{3}x-2p^{4}-p^{2}x-p^{3}+2p^{2}kx-kpx-kp^{3}x)$. Therefore
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}$
$\qquad=\frac{f_{1}(p,k,x)}{p^{4}(2p^{2}+p)(p^{5}x+2p^{4}x-2p^{3}x-2p^{4}-p^{2}x-p^{3}+2p^{2}kx-kpx-kp^{3}x)}$
$\qquad:=\frac{f_{1}(p,k,x)}{g_{1}(p,k,x)},$
where
$f_{1}(p,k,x)$ $=p^{13}x^{3}-4p^{11}x^{3}-2p^{12}x^{2}+2p^{10}x^{2}+2p^{10}x^{3}-2p^{8}x^{3}+2p^{11}x^{2}-2p^{9}x^{2}$
$-kp^{6}x^{3}-10kp^{5}x^{3}+4kp^{4}x^{3}+2kp^{3}x^{3}+8kp^{6}x^{2}+4kp^{5}x^{2}-4kp^{9}x^{3}-2kp^{8}x^{3}$
$+11kp^{7}x^{3}+4kp^{9}x^{2}+14kp^{8}x^{2}+3p^{9}x^{3}-kp^{2}x^{3}-2kp^{4}x^{2}-24kp^{7}x^{2}$
$-4kp^{10}x^{2}+5k^{2}p^{4}x^{3}-4k^{2}p^{3}x^{3}+12k^{2}p^{6}x^{2}-8k^{2}p^{5}x^{2}+4k^{2}p^{7}x^{3}$
$-5k^{2}p^{6}x^{3}-8k^{2}p^{7}x^{2}+k^{2}p^{2}x^{3}+2k^{2}p^{4}x^{2}+kp^{11}x^{3}-k^{2}p^{8}x^{3}+2k^{2}p^{8}x^{2}.$
On factorization, we have
$f_{1}(p,k,x)$ $=p^{2}x^{2}(p-1)^{2}\left(p^{6}x(p^{3}-p-2)+p^{7}(px-2p-2)+p^{4}x(p^{4}-k^{2})\right.$
$\quad\quad\quad\quad\quad\quad\quad\left.+2k^{2}p^{2}(p^{2}+1)+kx(k-1)+kp^{4}(x(p^{3}-p-6)-4p+10)\right.$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.+kp^{2}(5x-2)+2k^{2}p(p^{2}x-x-2p^{2})+2kp^{6}(x-2)\right).$
For all prime $p\geq 2,1\leq k\leq p$ and $x\geq p^{2}$, we have $p^{3}-p-2>0,px-2p-2>0,p^{4}-k^{2}>0,p^{2}x-x-2p^{2}>0$ and $x(p^{3}-p-6)>4p-10$. As such, $f_{1}(p,k,x)>0$. Also, $g_{1}(p,k,x)=\frac{2p^{10}}{x^{2}(p-1)^{2}}\times|e(\mathcal{CCC}(G))|\times|v(\mathcal{CCC}(G))|>0$. Therefore, $\frac{f_{1}(p,k,x)}{g_{1}(p,k,x)}>0$.
If $\mathcal{CCC}(G)=(kp+1)K_{n_{1}}\sqcup(p+1-k)K_{n_{2}}$ then, by Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G)$ $=(kp+1)\frac{x(p-1)}{p^{2}}\left(\frac{x(p-1)}{p^{2}}-1\right)^{2}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+(p+1-k)\frac{x(p-1)}{p}\left(\frac{x(p-1)}{p}-1\right)^{2}$
$=\frac{x(p-1)}{p^{6}}\left(p^{6}x^{2}+p^{6}+p^{5}-2kp^{2}x^{2}+kpx^{2}-4kp^{4}x+2kp^{3}x+p^{2}x^{2}\right.$
$\quad\quad\quad\quad\left.-2px^{2}+x^{2}-2p^{3}x+2p^{2}x+p^{4}-p^{5}x^{2}-p^{4}x^{2}-2p^{6}x+p^{3}x^{2}\right.$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.+2p^{4}x-kp^{5}x^{2}+2kp^{4}x^{2}+2kp^{5}x\right)\text{ \quad and }$
$M_{2}(\mathcal{CCC}(G))$ $=(kp+1)\frac{x(p-1)}{p^{2}}\frac{\left(\frac{x(p-1)}{p^{2}}-1\right)^{3}}{2}+(p+1-k)\frac{x(p-1)}{p}\frac{\left(\frac{x(p-1)}{p}-1\right)^{3}}{2}$
$=\frac{x(p-1)}{2p^{8}}\left(2kp^{4}x^{3}-3kp^{3}x^{3}+3kp^{2}x^{3}-kpx^{3}+6kp^{4}x^{2}-3kp^{3}x^{2}+p^{3}x^{3}\right.$
$\quad\quad\quad\quad\quad\left.+6kp^{6}x-3kp^{5}x-3p^{2}x^{3}+3px^{3}-x^{3}-3p^{4}x^{2}+6p^{3}x^{2}-p^{6}\right.$
$\quad\quad\quad\quad\quad\quad\left.-3p^{2}x^{2}+3p^{5}x-3p^{4}x-2p^{7}x^{3}+2p^{5}x^{3}-3p^{8}x^{2}+3p^{7}x^{2}\right.$
$\quad\quad\quad\quad\quad\quad\left.+p^{8}x^{3}+3p^{8}x-3p^{5}x^{2}-p^{8}-p^{4}x^{3}+3p^{6}x^{2}-3p^{6}x-p^{7}\right.$
$\quad\quad\quad\quad\quad\quad\left.-kp^{7}x^{3}+3kp^{6}x^{3}-3kp^{5}x^{3}+3kp^{7}x^{2}-6kp^{6}x^{2}-3kp^{7}x\right).$
Also, $|v(\mathcal{CCC}(G))|=(kp+1)\frac{x(p-1)}{p^{2}}+(p+1-k)\frac{x(p-1)}{p}=\frac{x(p-1)}{p^{2}}(p^{2}+p+1)$ and $|e(\mathcal{CCC}(G))|=(kp+1)\binom{\frac{x(p-1)}{p^{2}}}{2}+(p+1-k)\binom{\frac{x(p-1)}{p}}{2}=\frac{x(p-1)}{2p^{4}}(2kp^{2}x-kpx+px-x-p^{2}+p^{4}x+p^{4}-p^{2}x-p^{3}-kp^{3}x)$. Therefore
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}$
$\!\!\!\!\!=\frac{f_{2}(p,k,x)}{p^{4}(p^{2}+p+1)(2kp^{2}x-kpx+px-x-p^{2}+p^{4}x+p^{4}-p^{2}x-p^{3}-kp^{3}x)}:=\frac{f_{2}(p,k,x)}{g_{2}(p,k,x)},$
where
$f_{2}(p,k,x)$ $=kp^{9}x^{3}-k^{2}p^{8}x^{3}-3kp^{8}x^{3}+p^{8}x^{3}+4k^{2}p^{7}x^{3}-3p^{7}x^{3}-5k^{2}p^{6}x^{3}+9kp^{6}x^{3}$
$\quad+p^{6}x^{3}-10kp^{5}x^{3}+5p^{5}x^{3}+5k^{2}p^{4}x^{3}-kp^{4}x^{3}-5p^{4}x^{3}-4k^{2}p^{3}x^{3}+8kp^{3}x^{3}$
$\quad-p^{3}x^{3}+k^{2}p^{2}x^{3}-5kp^{2}x^{3}+3p^{2}x^{3}+kpx^{3}-px^{3}-2kp^{9}x^{2}+2k^{2}p^{8}x^{2}$
$\quad+6kp^{8}x^{2}-2p^{8}x^{2}-8k^{2}p^{7}x^{2}-2kp^{7}x^{2}+6p^{7}x^{2}+12k^{2}p^{6}x^{2}-12kp^{6}x^{2}$
$\quad-4p^{6}x^{2}-8k^{2}p^{5}x^{2}+18kp^{5}x^{2}-4p^{5}x^{2}+2k^{2}p^{4}x^{2}-10kp^{4}x^{2}+6p^{4}x^{2}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+2kp^{3}x^{2}-2p^{3}x^{2}.$
On factorization, we have $f_{2}(p,k,x)=px^{2}(p-1)^{4}(p+1-k)(kp+1)(p^{2}x-x-2p^{2})$. For all $p\geq 2,1\leq k\leq p$ and $x\geq p^{2}$, we have $p+1-k>0$ and $x(p^{2}-1)>2p^{2}$. As such, $f_{2}(p,k,x)>0$. Also, $g_{2}(p,k,x)=\frac{2p^{10}}{x^{2}(p-1)^{2}}\times|e(\mathcal{CCC}(G))|\times|v(\mathcal{CCC}(G))|>0$. Therefore, $\frac{f_{2}(p,k,x)}{g_{2}(p,k,x)}>0$.
If $\mathcal{CCC}(G)=K_{m}\sqcup pK_{n_{2}}$ then, by Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))$ $=\frac{x(p^{2}-1)}{p}\left(\frac{x(p^{2}-1)}{p}-1\right)^{2}+p\frac{x(p-1)}{p}\left(\frac{x(p-1)}{p}-1\right)^{2}$
$=\frac{x(p-1)}{p^{3}}\left(p^{5}x^{2}-p^{3}x^{2}+2px^{2}-2p^{4}x+2p^{3}+p^{4}x^{2}-4p^{2}x^{2}+x^{2}\right.$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.+2px-4p^{3}x+4p^{2}x+p^{2}\right)$
and
$M_{2}$ $(\mathcal{CCC}(G))=\frac{x(p^{2}-1)}{p}\frac{\left(\frac{x(p^{2}-1)}{p}-1\right)^{3}}{2}+p\frac{x(p-1)}{p}\frac{\left(\frac{x(p-1)}{p}-1\right)^{3}}{2}$
$=\frac{x(p-1)}{2p^{4}}\left(p^{7}x^{3}-3p^{5}x^{3}-2px^{3}-3p^{6}x^{2}+3p^{4}x^{2}-6p^{2}x^{2}+3p^{5}x-6p^{3}x\right.$
$\left.-2p^{4}+p^{6}x^{3}-2p^{4}x^{3}+6p^{2}x^{3}-x^{3}-3p^{5}x^{2}+12p^{3}x^{2}-3px^{2}+6p^{4}x-3p^{2}x-p^{3}\right.$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.+6p^{4}x-3p^{2}x-p^{3}\right).$
Also, $|v(\mathcal{CCC}(G))|=\frac{x(p^{2}-1)}{p}+p\frac{x(p-1)}{p}=\frac{x(p-1)}{p}(2p+1)$ and $|e(\mathcal{CCC}(G))|=\binom{\frac{x(p^{2}-1)}{p}}{2}+p\binom{\frac{x(p-1)}{p}}{2}=\frac{x(p-1)}{2p^{2}}(p^{3}x-2px+2p^{2}x-2p^{2}-x-p)$. Therefore, for all $p\geq 2$ and $x\geq p^{2}$, we have
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}$
$\qquad\qquad=\frac{p^{8}x^{3}-4p^{6}x^{3}+2p^{5}x^{3}+3p^{4}x^{3}-2p^{3}x^{3}-2p^{7}x^{2}+2p^{6}x^{2}+2p^{5}x^{2}-2p^{4}x^{2}}{p^{2}(2p+1)(p^{3}x-2px+2p^{2}x-2p^{2}-x-p)}$
$\qquad\qquad=\frac{p^{6}x^{3}(p^{2}-4)+2p^{5}x^{2}(x-p^{2})+p^{3}x^{3}(3p-2)+2p^{6}x^{2}+2p^{4}x^{2}(p-1)}{p^{2}(2p+1)\{x(p^{3}-2p-1)+p(2px-2p-1)\}}>0.$
If $\mathcal{CCC}(G)=(p^{2}+p+1)K_{n_{1}}$ then the result follows from Theorem 2.14. If $\mathcal{CCC}(G)=K_{n_{1}}\sqcup(p+1)K_{n_{2}}$ then, by Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))$ $=\frac{x(p-1)}{p^{2}}\left(\frac{x(p-1)}{p^{2}}-1\right)^{2}+(p+1)\frac{x(p-1)}{p}\left(\frac{x(p-1)}{p}-1\right)^{2}$
$=\frac{x(p-1)}{p^{6}}\left(p^{6}x^{2}-p^{5}x^{2}-p^{4}x^{2}+p^{3}x^{2}-2p^{6}x+p^{6}+2p^{4}x+p^{5}+p^{4}\right.$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.+p^{2}x^{2}-2px^{2}+x^{2}-2p^{3}x+2p^{2}x\right)$
and
$M_{2}(\mathcal{CCC}(G))$ $=\frac{x(p-1)}{p^{2}}\frac{\left(\frac{x(p-1)}{p^{2}}-1\right)^{3}}{2}+(p+1)\frac{x(p-1)}{p}\frac{\left(\frac{x(p-1)}{p}-1\right)^{3}}{2}$
$=\frac{x(p-1)}{2p^{8}}\left(p^{8}x^{3}-2p^{7}x^{3}+2p^{5}x^{3}-p^{4}x^{3}-3p^{8}x^{2}+3p^{7}x^{2}+3p^{6}x^{2}\right.$
$\quad\quad\quad\quad\quad\left.-3p^{5}x^{2}+3p^{8}x-3p^{6}x+3p^{5}x-3p^{4}x-p^{8}-p^{7}-p^{6}+p^{3}x^{3}\right.$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.-3p^{2}x^{3}+3px^{3}-x^{3}-3p^{4}x^{2}+6p^{3}x^{2}-3p^{2}x^{2}\right).$
Also, $|v(\mathcal{CCC}(G))|=\frac{x(p-1)}{p^{2}}+(p+1)\frac{x(p-1)}{p}=\frac{x(p-1)}{p^{2}}(p^{2}+p+1)$ and $|e(\mathcal{CCC}(G))|=\binom{\frac{x(p-1)}{p^{2}}}{2}+(p+1)\binom{\frac{x(p-1)}{p}}{2}=\frac{x(p-1)}{2p^{4}}(p^{4}x+px-p^{2}x-p^{4}-p^{3}-p^{2}-x)$. Therefore, $\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{f_{3}(p,x)}{p^{4}(p^{2}+p+1)(p^{4}x+px-p^{2}x-p^{4}-p^{3}-p^{2}-x)}:=\frac{f_{3}(p,x)}{g_{3}(p,x)}$, where
$f_{3}(p,x)$ $=p^{8}x^{3}-3p^{7}x^{3}+p^{6}x^{3}+5p^{5}x^{3}-5p^{4}x^{3}-p^{3}x^{3}+3p^{2}x^{3}-px^{3}-2p^{8}x^{2}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+6p^{7}x^{2}-4p^{6}x^{2}-4p^{5}x^{2}+6p^{4}x^{2}-2p^{3}x^{2}.$
On factorization, we have $f_{3}(p,x)=px^{2}(p+1)(p-1)^{4}(p^{2}x-x-2p^{2})$. For all prime $p\geq 2$ and $x\geq p^{2}$, we have $x(p^{2}-1)>2p^{2}$. As such, $f_{3}(p,x)>0$ and $g_{3}(p,x)=\frac{2p^{10}}{x^{2}(p-1)^{2}}\times|e(\mathcal{CCC}(G))|\times|v(\mathcal{CCC}(G))|>0$. Therefore, $\frac{f_{3}(p,x)}{g_{3}(p,x)}>0$.
In all the cases, we have $\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}>\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}$. Hence, the result follows. ∎
Corollary 2.17.
Let $G$ be a non-abelian group of order $p^{n}$ such that $\frac{G}{Z(G)}$ is non-abelian and $|Z(G)|=p^{n-3}$, where $p$ is a prime and $n\geq 4$. Then $\mathcal{CCC}(G)$ satisfies Hansen-Vukičević conjecture.
Proof.
Clearly $|\frac{G}{Z(G)}|=p^{3}$ and so the result follows from Theorem 2.16. ∎
Corollary 2.18.
If $G$ is a non-abelian group of order $p^{4}$ then $\mathcal{CCC}(G)$ satisfies Hansen-Vukičević conjecture.
Proof.
We have $|G|=p^{4}$, $p$ is a prime, then $|Z(G)|$ must be $p$ or $p^{2}$. If $|Z(G)|=p^{2}$ then the result follows from Corollary 2.10. Otherwise, if $|Z(G)|=p$ then the result follows from Corollary 2.17. ∎
3. Concluding remark
In this paper we have shown that $\mathcal{CCC}(G)$ satisfies Hansen-Vukičević conjecture if
  • $G$ is isomorphic to $D_{2m},Q_{4m},SD_{8m},V_{8m},U(n,m)$ and $G(p,m,n)$.
  • $\frac{G}{Z(G)}$ is isomorphic to $D_{2m},\mathbb{Z}_{p}\times\mathbb{Z}_{p}$ and Frobenious group of order $pq$ and $p^{2}q$ for any two primes $p$ and $q$.
  • $|\frac{G}{Z(G)}|=p^{3}$ for any prime $p.$
As a consequence it was also seen that if $G$ is any non-abelian group of order $p^{3}$ or $p^{4}$, where $p$ is any prime, then $\mathcal{CCC}(G)$ satisfies Hansen-Vukičević conjecture. Using Theorem 2.8, Theorem 2.9 and the characterizations of $4$-centralizer (“$G$ is a $4$-centralizer finite group if and only if $\frac{G}{Z(G)}\cong\mathbb{Z}_{2}\times\mathbb{Z}_{2}$[3, Theorem 2]), $5$-centralizer (“$G$ is a $5$-centralizer finite group if and only if $\frac{G}{Z(G)}\cong\mathbb{Z}_{3}\times\mathbb{Z}_{3}$ or $D_{2\times 3}$[3, Theorem 4]) finite groups we have the following result.
Theorem 3.1.
If $G$ is a $4$-centralizer or $5$-centralizer finite group then $\mathcal{CCC}(G)$ satisfies Hansen-Vukičević conjecture.
Using Theorem 2.9 and the characterization of $(p+2)$-centralizer finite $p$-group (“If $G$ is a finite $p$-group then it is $(p+2)$-centralizer if and only if $\frac{G}{Z(G)}\cong\mathbb{Z}_{p}\times\mathbb{Z}_{p}$[2, Lemma 2.7]) we have the following result.
Theorem 3.2.
If $G$ is a $(p+2)$-centralizer finite $p$-group then $\mathcal{CCC}(G)$ satisfies Hansen-Vukičević conjecture.
If $\{g_{1},g_{2},\dots,g_{m}\}$ is a set of pairwise non-commuting elements of $G$ having maximal size then $G$ is a $4$-centralizer group if $m=3$ and a $5$-centralizer group if $m=4$ (see [1, Lemma 2.4]). Therefore, if $m=3$ and $4$ then $\mathcal{CCC}(G)$ graphs of such groups satisfy Hansen-Vukičević conjecture.
Let $p$ be the smallest prime divisor of $|G|$. Then the fact that “commuting probability of $G$, $\Pr(G)=\frac{p^{2}+p-1}{p^{3}}$ if and only if $\frac{G}{Z(G)}\cong\mathbb{Z}_{p}\times\mathbb{Z}_{p}$[20, Theorem 3] along with Theorem 2.9 give the following result.
Theorem 3.3.
If $p$ is the smallest prime divisor of $|G|$ and $\Pr(G)=\frac{p^{2}+p-1}{p^{3}}$ then $\mathcal{CCC}(G)$ satisfies Hansen-Vukičević conjecture.
Further, if $\Pr(G)=\frac{5}{14},\frac{2}{5},\frac{11}{27},\frac{1}{2}$ or $\frac{5}{8}$ then $\frac{G}{Z(G)}\cong D_{2\times 7},D_{2\times 5},D_{2\times 4},D_{2\times 3}$ or $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ (see [24, pp. 246] and [21, pp. 451]). Therefore, the following result follows in view of Theorem 2.8 and Theorem 2.9.
Theorem 3.4.
If $\Pr(G)=\frac{5}{14},\frac{2}{5},\frac{11}{27},\frac{1}{2}$ or $\frac{5}{8}$ then $\mathcal{CCC}(G)$ satisfies Hansen-Vukičević conjecture.
In our study we have not found any group $G$ such that $\mathcal{CCC}(G)$ does not satisfy Hansen-Vukičević conjecture. We conclude this paper with the following question.
Question 3.5.
Is there any finite group $G$ such that $\mathcal{CCC}(G)$ does not satisfy Hansen-Vukičević conjecture?
Acknowledgement. The first author is thankful to Council of Scientific and Industrial Research for the fellowship (File No. 09/0796(16521)/2023-EMR-I).
Declarations
  • Funding: No funding was received by the authors.
  • Conflict of interest: The authors declare that they have no conflict of interest.
  • Availability of data and materials: No data was used in the preparation of this manuscript.